### v1

# MsgABD: EveryShipout initializing macros

Abstract
Message Passing Neural Networks (MPNNs) are a class of Graph Neural Networks (GNNs) that leverage the graph topology to propagate messages across increasingly larger neighborhoods. The message-passing scheme leads to two distinct challenges: over-smoothing and over-squashing. While several algorithms, e.g. DropEdge and its variants – DropNode, DropAgg and DropGNN – have successfully addressed the over-smoothing problem, their impact on over-squashing remains largely unexplored. This represents a critical gap in the literature as failure to mitigate over-squashing would make these methods unsuitable for long-range tasks. In this work, we take the first step towards closing this gap by studying the aforementioned algorithms in the context of over-squashing. We present novel theoretical results that characterize the negative effects of DropEdge on sensitivity between distant nodes, suggesting its unsuitability for long-range tasks. Our findings are easily extended to its variants, allowing us to build a comprehensive understanding of how they affect over-squashing. We evaluate these methods using real-world datasets, demonstrating their detrimental effects. Specifically, we show that while DropEdge-variants improve test-time performance in short-range tasks, they deteriorate performance in long-range ones. Our theory explains these results as follows: random edge-dropping lowers the effective receptive field of GNNs, which although beneficial for short-range tasks, misaligns the models on long-range ones. This forces the models to overfit to short-range artefacts in the training set, resulting in poor generalization. Our conclusions highlight the need to re-evaluate various methods designed for training deep GNNs, with a renewed focus on modelling long-range interactions.
1 Introduction
Graph-structured data is ubiquitous – it is found in social media platforms, online retail platforms, molecular structures, transportation networks, and even computer systems. Graph neural networks (GNNs) (Scarselli et al., 2009; Li et al., 2016) are powerful neural models developed for modelling graph-structured data, and have found applications in several real-world scenarios (Gao et al., 2018; You et al., 2020c; a; 2022; b; Ying et al., 2018; Zheng et al., 2022; Monti et al., 2017; Zitnik & Leskovec, 2017; Wale & Karypis, 2006). A popular class of GNNs, called message-passing neural networks (MPNNs) (Gilmer et al., 2017), recursively process neighborhood information using message-passing layers. These layers are stacked to allow each node to aggregate information from increasingly larger neighborhoods, akin to how convolutional neural networks (CNNs) learn hierarchical features for images (LeCun et al., 1989). However, unlike in image-based deep learning, where ultra-deep CNN architectures have led to performance breakthroughs (Szegedy et al., 2015; He et al., 2016), shallow GNNs often outperform deeper models on many graph learning tasks (Zhou et al., 2021b). This is because deep GNNs suffer from unique issues like over-smoothing (Oono & Suzuki, 2020) and over-squashing (Alon & Yahav, 2021), which makes training them notoriously difficult.
Over-smoothing refers to the problem of node representations becoming too similar as they are recursively processed. This is undesirable since it limits the GNN from effectively utilizing the information in the input features. The problem has garnered significant attention from the research community, resulting in a suite of algorithms designed to address it Rusch et al. (2023). Amongst these methods are a collection of random edge-dropping algorithms, including DropEdge (Rong et al., 2020) and its variants – DropNode (Feng et al., 2020), DropAgg Jiang et al. (2023) and DropGNN (Papp et al., 2021) – which act as message-passing reducers.
The other issue specific to GNNs is over-squashing. In certain graph structures, neighborhood size grows exponentially with distance from the source (Chen et al., 2018b), causing information to be lost as it passes through graph bottlenecks (Alon & Yahav, 2021). This limits MPNNs’ ability to enable communication between distant nodes, which is crucial for good performance on long-range tasks. To alleviate over-squashing, several graph-rewiring techniques have been proposed, which aim to improve graph connectivity by adding edges in a strategic manner [^1] Sometimes, along with removal of some edges to preserve statistical properties of the original topology. (Karhadkar et al., 2023; Deac et al., 2022; Black et al., 2023; Nguyen et al., 2023; Alon & Yahav, 2021). In contrast, the DropEdge-variants only remove edges, which should, in principle, amplify over-squashing levels.
The empirical evidence in support of methods designed for training deep GNNs has been majorly collected on short-range tasks. That is, it simply suggests that these methods prevent loss of local information, but it remains inconclusive if they facilitate capturing long-range interactions (LRIs). Of course, on long-range tasks, deeper GNNs are useless if they cannot capture LRIs. This is especially a concern for DropEdge-variants since evidence suggests that alleviating over-smoothing with graph rewiring could exacerbate over-squashing (Nguyen et al., 2023; Giraldo et al., 2023).
Theoretical Contributions. In this work, we uncover the effects of random edge-dropping algorithms on over-squashing in MPNNs. By explicitly computing the expected sensitivity of the node representations to the node features (Topping et al., 2022) (inversely related to over-squashing) in a linear GCN Kipf & Welling (2017), we show that these methods provably reduce the effective receptive field of the model. Precisely speaking, the rate at which sensitivity between distant nodes decays is polynomial w.r.t. the dropping probability. Finally, we extend the existing theoretical results on sensitivity in nonlinear MPNNs Black et al. (2023); Di Giovanni et al. (2023); Xu et al. (2018) to the random edge-dropping setting, again showing that these algorithms exacerbate the over-squashing problem.
Experimental Results. We evaluate the DropEdge-variants on node classification tasks using GCN and GAT Veličković et al. (2018) architectures. Specifically, we assess their performance on homophilic datasets – Cora McCallum et al. (2000) and CiteSeer Giles et al. (1998) – which represent short-range tasks, and heterophilic datasets – Chameleon, Squirrel, TwitchDE Rozemberczki et al. (2021) – which correspond to long-range tasks. Our results indicate an increasing trend in test accuracy for homophilic datasets and a declining trend for heterophilic datasets as the dropping probability increases. Accordingly, we hypothesize that edge-dropping algorithms improve performance on short-range tasks, as has been reported earlier, by reducing the receptive field of the GNN and increasing model-dataset alignment, However, for long-range tasks, they decrease the model-dataset alignment, resulting in poor generalization.
2 Background
Consider a directed graph $\mathcal{G}=\left(\mathcal{V},\mathcal{E}\right)$, with $\mathcal{V}=\left[N\right]\coloneqq\left\{1,\ldots,N\right\}$ denoting the node set and $\mathcal{E}\subset\mathcal{V}\times\mathcal{V}$ the edge set; $\left(j\to i\right)\in\mathcal{E}$ if there’s an edge from node $j$ to node $i$. Let $\bm{A}\in\left\{0,1\right\}^{N\times N}$ denote its adjacency matrix, such that $\bm{A}_{ij}=1$ if and only if $\left(j\to i\right)\in\mathcal{E}$, and let $\bm{D}\coloneqq\text{diag}\left(\bm{A}\mathbf{1}_{N}\right)$ denote the in-degree matrix. The geodesic distance, $d_{\mathcal{G}}\left(j,i\right)$, from node $j$ to node $i$ is the length of the shortest path starting at node $j$ and ending at node $i$. Accordingly, the $\ell$-hop neighborhood of a node $i$ can be defined as the set of nodes that can reach it in exactly $\ell\in\mathbb{N}_{0}$ steps, $\mathbb{S}^{\left(\ell\right)}\left(i\right)=\left\{j\in\mathcal{V}:d_{\mathcal{G}}\left(j,i\right)=\ell\right\}$.
2.1 Graph Neural Networks
Graph neural networks (GNNs) operate on inputs of the form $\left(\mathcal{G},\bm{X}\right)$, where $\mathcal{G}$ encodes the graph topology and $\bm{X}\in^{N\times H^{\left(0\right)}}$ collects the node features [^2] To keep things simple, we will ignore edge features.. Message-passing neural networks (MPNNs) Gilmer et al. (2017) are a special class of GNNs which recursively aggregate information from the 1-hop neighborhood of each node using message-passing layers. An L-layer MPNN is given as
$\begin{split}&\bm{z}_{i}^{\left(\ell\right)}=\mathsf{Upd}^{\left(\ell\right)}\left(\bm{z}_{i}^{\left(\ell-1\right)},\mathsf{Agg}^{\left(\ell\right)}\left(\bm{z}_{i}^{\left(\ell-1\right)},\left\{\bm{z}_{j}^{\left(\ell-1\right)}:j\in\mathbb{S}^{\left(1\right)}\left(i\right)\right\}\right)\right),\quad\forall\ell\in\left[L\right]\\ &\mathsf{MPNN}_{\theta}\left(\mathcal{G},\bm{X}\right)=\left\{\mathsf{Out}\left(\bm{z}_{i}^{\left(L\right)}\right):i\in\mathcal{V}\right\}\end{split}$ (2.1)
where $\bm{Z}^{\left(0\right)}=\bm{X}$, $\mathsf{Agg}^{\left(\ell\right)}$ denotes the aggregation functions, $\mathsf{Upd}^{\left(\ell\right)}$ the update functions, and $\mathsf{Out}$ the readout function. Since the final representation of node $i$ is a function of the input features of nodes at most L-hops away from it, its receptive field is given by $\mathbb{B}^{\left(L\right)}\left(i\right)\coloneqq\left\{j\in\mathcal{V}:d_{\mathcal{G}}\left(j,i\right)\leq L\right\}$.
For example, a GCN Kipf & Welling (2017) updates node representations as the weighted sum of its neighbors’ representations:
$\bm{Z}^{\left(\ell\right)}=\sigma\left(\hat{\bm{A}}\bm{Z}^{\left(\ell-1\right)}\bm{W}^{\left(\ell\right)}\right)$ (2.2)
where $\sigma$ is a point-wise nonlinearity, e.g. ReLU, the propagation matrix, $\hat{\bm{A}}$, is a graph shift operator, i.e. $\hat{\bm{A}}_{ij}\neq 0$ if and only if $\left(j\to i\right)\in\mathcal{E}$ or $i=j$, and $\bm{W}^{\left(\ell\right)}\in^{H^{\left(\ell-1\right)}\times H^{\left(\ell\right)}}$ is a weight matrix. The original choice for $\hat{\bm{A}}$ was the symmetrically normalized adjacency matrix $\hat{\bm{A}}^{\text{sym}}\coloneqq\widetilde{\bm{D}}^{-1/2}\widetilde{\bm{A}}\widetilde{\bm{D}}^{-1/2}$ Kipf & Welling (2017), where $\widetilde{\bm{A}}=\bm{A}+\bm{I}_{N}$ and $\widetilde{\bm{D}}=\text{diag}(\widetilde{\bm{A}}\mathbf{1}_{N})$. However, several influential works have also used the asymmetrically normalized adjacency, $\hat{\bm{A}}^{\text{asym}}\coloneqq\widetilde{\bm{D}}^{-1}\widetilde{\bm{A}}$ Hamilton et al. (2017); Schlichtkrull et al. (2017); Li et al. (2018).
2.2 DropEdge
DropEdge is a random data augmentation technique that works by sampling a subgraph of the original input graph in each layer, and uses that for message passing Rong et al. (2020):
$\begin{split}\mathbf{M}^{\left(\ell\right)}&\sim\left\{\text{Bern}\left(1-q\right)\right\}^{N\times N}\\ \widetilde{\bm{A}}^{\left(\ell\right)}&=\mathbf{M}^{\left(\ell\right)}\circ\bm{A}+\bm{I}_{N}\end{split}$ (2.3)
Several variants of DropEdge have also been proposed, forming a family of random edge-dropping algorithms for tackling the over-smoothing problem. For example, DropNode (Feng et al., 2020) independently samples nodes and sets their representations to 00, followed by rescaling to make the feature matrix unbiased. This is equivalent to setting the corresponding columns of the propagation matrix to 00. In a similar vein, DropAgg (Jiang et al., 2023) samples nodes that don’t aggregate messages from their neighbors. This is equivalent to dropping the corresponding rows of the adjacency matrix. Combining these two approaches, DropGNN (Papp et al., 2021) samples nodes which neither propagate nor aggregate messages in a given message-passing step. These algorithms alleviate over-smoothing by reducing the number of messages being propagated in the graph, thereby slowing down the convergence of node representations.
2.3 Over-squashing
Over-squashing refers to the problem of information from exponentially growing neighborhoods Chen et al. (2018a) being squashed into finite-sized node representations Alon & Yahav (2021). This results in a loss of information as it is propagated over long distances, disallowing MPNNs from capturing long-range interactions (LRIs) and limiting their applications to short-range tasks. Topping et al. (2022) formally characterized over-squashing in terms of the Jacobian of the node-level representations w.r.t. the input features: $\|\partial\bm{z}^{\left(L\right)}_{i}/\partial\bm{x}_{j}\|_{1}$. Accordingly, over-squashing can be understood as low sensitivity between distant nodes, i.e. small perturbations in a node’s features don’t effect other distant nodes’ representations.
Several works have linked over-squashing in an MPNN with topological properties like Cheeger’s constant Giraldo et al. (2023); Karhadkar et al. (2023), curvature of edges Topping et al. (2022); Nguyen et al. (2023); Liu et al. (2023), effective resistance between nodes Black et al. (2023); Arnaiz-Rodríguez et al. (2022) and the expected commute time between them Di Giovanni et al. (2023); Giovanni et al. (2024). These results have inspired the design of several graph rewiring techniques that strategically add edges to improve the connectivity in the graph, thereby alleviating over-squashing.
3 Sensitivity Analysis
In this section, we perform a theoretical analysis of the expectation – w.r.t. random edge masks – of sensitivity of node representations. This will allow us to predict how DropEdge-variants affect communication between nodes at various distances, which is relevant for predicting their suitability towards learning LRIs.
3.1 Linear GCNs
We start our analysis with linear GCNs, and treat more general MPNN architectures in the following subsection. In this model, the final node representations can be summarised as
$\bm{Z}^{\left(L\right)}=\left(\prod_{\ell=1}^{L}\hat{\bm{A}}^{\left(\ell\right)}\right)\bm{X}\bm{W}\in^{N\times H^{\left(L\right)}}$ (3.1)
where $\bm{W}\coloneqq\prod_{\ell=1}^{L}\bm{W}^{\left(\ell\right)}\in^{H^{\left(0\right)}\times H^{\left(L\right)}}$. Using the i.i.d. assumption on the distribution of edge masks in each layer, the expected sensitivity of node $i$ to node $j$ can be shown to be
$\mathbb{E}_{\mathbf{M}^{\left(1\right)},\ldots,\mathbf{M}^{\left(L\right)}}\left[\left\|\frac{\partial\bm{z}^{\left(L\right)}_{i}}{\partial\bm{x}_{j}}\right\|_{1}\right]=\left(\mathbb{E}\left[\hat{\bm{A}}\right]^{L}\right)_{ij}\left\|\bm{W}\right\|_{1}$ (3.2)
To keep things simple, we will ignore the effect of DropEdge-variants on the optimization trajectory. Accordingly, it is sufficient to study $\mathbb{E}[\hat{\bm{A}}]$ in order to predict their effect on over-squashing. To maintain analytical tractability, we assume the use of an asymmetrically normalized adjacency matrix for message-passing, $\hat{\bm{A}}=\hat{\bm{A}}^{\text{asym}}$.
Lemma 3.1.
The expected propagation matrix under DropEdge is given as:
$\begin{split}\dot{\bm{P}}_{ii}&\coloneqq\mathbb{E}_{\mathsf{DE}}\left[\hat{\bm{A}}_{ii}\right]=\frac{1-q^{d_{i}+1}}{\left(1-q\right)\left(d_{i}+1\right)}\\ \dot{\bm{P}}_{ij}&\coloneqq\mathbb{E}_{\mathsf{DE}}\left[\hat{\bm{A}}_{ij}\right]=\frac{1}{d_{i}}\left(1-\frac{1-q^{d_{i}+1}}{\left(1-q\right)\left(d_{i}+1\right)}\right)\end{split}$ (3.3)
where $q\in[0,1)$ is the dropping probability.
See Section B.1 for a proof.
Other Variants. We will similarly derive the expected propagation matrix for other random edge-dropping algorithms. First off, DropNode (Feng et al., 2020) samples nodes and drops corresponding columns from the aggregation matrix directly, followed by rescaling of its entries:
$\begin{split}\mathbb{E}_{\mathsf{DN}}\left[\frac{1}{1-q}\hat{\bm{A}}\right]=\frac{1}{1-q}\times\left(1-q\right)\hat{\bm{A}}=\hat{\bm{A}}\end{split}$ (3.4)
That is, the expected propagation matrix is the same as in a NoDrop model ($q=0$).
Nodes sampled by DropAgg Jiang et al. (2023) don’t aggregate messages. Therefore, if $\hat{\bm{A}}=\hat{\bm{A}}^{\text{asym}}$, then the expected propagation matrix is given by
$\begin{split}\mathbb{E}_{\mathsf{DA}}\left[\hat{\bm{A}}_{ii}\right]&=q+\frac{1-q}{d_{i}+1}=\frac{1+d_{i}q}{d_{i}+1}>\mathbb{E}_{\mathsf{DE}}\left[\hat{\bm{A}}_{ii}\right]\\ \mathbb{E}_{\mathsf{DA}}\left[\hat{\bm{A}}_{ij}\right]&=\frac{1}{d_{i}}\left(1-\mathbb{E}_{\mathsf{DA}}\left[\hat{\bm{A}}_{ii}\right]\right)<\mathbb{E}_{\mathsf{DE}}\left[\hat{\bm{A}}_{ij}\right]\end{split}$ (3.5)
Finally, DropGNN (Papp et al., 2021) samples nodes which neither propagate nor aggregate messages. From any node’s perspective, if it is not sampled, then its aggregation weights are computed as for DropEdge:
$\begin{split}\mathbb{E}_{\mathsf{DG}}\left[\hat{\bm{A}}_{ii}\right]&=q+\left(1-q\right)\mathbb{E}_{\mathsf{DE}}\left[\hat{\bm{A}}_{ii}\right]=q+\frac{1-q^{d_{i}+1}}{d_{i}+1}>\mathbb{E}_{\mathsf{DA}}\left[\hat{\bm{A}}_{ii}\right]\\ \mathbb{E}_{\mathsf{DG}}\left[\hat{\bm{A}}_{ij}\right]&=\frac{1}{d_{i}}\left(1-\mathbb{E}_{\mathsf{DG}}\left[\hat{\bm{A}}_{ii}\right]\right)<\mathbb{E}_{\mathsf{DA}}\left[\hat{\bm{A}}_{ij}\right]\end{split}$ (3.6)
1-Layer Linear GCNs. $\forall q\in\left(0,1\right)$ we have
$\begin{split}\dot{\bm{P}}_{ii}&=\frac{1}{d_{i}+1}\sum_{k=0}^{d_{i}}q^{k}>\frac{1}{d_{i}+1}\\ \dot{\bm{P}}_{ij}&=\frac{1}{d_{i}}\left(1-\dot{\bm{P}}_{ii}\right)<\frac{1}{d_{i}+1}\end{split}$ (3.7)
where the right-hand sides of the two inequalities are the corresponding entries in the propagation matrix of a NoDrop model. Equation 3.3 to Equation 3.7 together imply the following result:
Lemma 3.2.
In a 1-layer linear GCN with $\hat{\bm{A}}=\hat{\bm{A}}^{\text{asym}}$, using DropEdge, DropAgg or DropGNN
  1. 1.
    increases the sensitivity of a node’s representations to its own input features, and
  2. 2.
    decreases the sensitivity to its neighbors’ features.
In other words, DropEdge-variants prevent a 1-layer GCN from fully utilizing neighborhood information when learning node representations. Ignoring the graph topology this way makes the model resemble an MLP, limiting its expressiveness and hindering its ability to model graph-data.
Figure 1: Entries of $\dot{\bm{P}}^{6}$, averaged over molecular graphs sampled from the Proteins dataset. Left: Sensitivity between nodes decays at a polynomial rate w.r.t. their distance. Middle: Similarly, it decays at a polynomial rate w.r.t. the DropEdge probability. Right: It decays more quickly with DropAgg than with DropEdge, and even more rapidly with DropGNN.
L-layer Linear GCNs. Unfortunately, we cannot draw similar conclusions in L-layer networks, for nodes at arbitrary distances. To see this, view $\dot{\bm{P}}$ as the transition matrix of a non-uniform random walk. This walk has higher self-transition ($i=j$) probabilities than in a uniform augmented random walk ($\bm{P}=\hat{\bm{A}}^{\text{asym}}$, $q=0$), but lower inter-node ($i\neq j$) transition probabilities. Note that $\dot{\bm{P}}^{L}$ and $\bm{P}^{L}$ store the L-step transition probabilities in the corresponding walks. Then, since the paths connecting the nodes $i\in\mathcal{V}$ and $j\in\mathbb{B}^{\left(L-1\right)}\left(i\right)$ may involve self-loops, $(\dot{\bm{P}}^{L})_{ij}$ may be lower or higher than $(\bm{P}^{L})_{ij}$. Therefore, we cannot conclude how sensitivity between nodes separated by at most $L-1$ hops changes. For nodes L-hops away, however, we can show that DropEdge always decreases the corresponding entry in $\dot{\bm{P}}^{L}$, reducing the effective reachability of GCNs. Using Equation 3.5 and Equation 3.6, we can show the same for DropAgg and DropGNN, respectively.
Theorem 3.1.
In an L-layer linear GCN with $\hat{\bm{A}}=\hat{\bm{A}}^{\text{asym}}$, using DropEdge, DropAgg or DropGNN decreases the sensitivity of a node $i\in\mathcal{V}$ to another node $j\in\mathbb{S}^{\left(L\right)}\left(i\right)$, thereby reducing its effective receptive field. Moreover, the sensitivity monotonically decreases as the dropping probability is increased.
See Section B.2 for a precise quantitative statement and the proof.
Nodes at Arbitrary Distances. Although no general statement could be made about the change in sensitivity between nodes up to $L-1$ hops away, we can analyze such pairs empirically. We sampled 100 molecular graphs from the Proteins dataset (Dobson & Doig, 2003), binned the node-pairs in each graph by the shortest distance between them, and then plotted the average of the corresponding entries in $\dot{\bm{P}}^{L}$, $L=6$.
The results are shown in Figure 1. In the left subfigure, we observe that the sensitivity between two nodes decays at a polynomial rate with increasing distance between them. Moreover, DropEdge increases the expected sensitivity between nodes close to each other (0-hop and 1-hop neighbors) in the original topology, but reduces it between nodes farther off. In the middle subfigure, we show how the average sensitivity at different distances changes with the DropEdge probability. Specifically, we observe that the decay in sensitivity to nodes at large distances is polynomial in the DropEdge probability, which suggests that the algorithm would not be suitable for capturing LRIs. Similar conclusions can be made with the symmetrically normalized propagation matrix (see Section D.1). Finally, in the right subfigure, we compare the DropEdge-variants by plotting the sensitivity at $d_{\mathcal{G}}\left(i,j\right)=L=6$. Although the analytical form of all the expected propagation matrices are available to us, we approximate them using Monte-Carlo sampling. As expected from Equation 3.5, we observe that DropAgg not only decreases the sensitivity of node representations to distant nodes, but does it to a greater extent than DropEdge. Similarly, Equation 3.6 suggested that DropGNN could be even more harmful than DropAgg, and this is validated by the empirical results. In fact, for $q>0.7$, nodes 6-hops away are mostly insensitive to each other.
3.2 Nonlinear MPNNs
While linear networks are useful in simplifying the theoretical analysis, they are often not practical. In this subsection, we will consider the upper bounds on sensitivity established in previous works, and extend them to the DropEdge setting.
ReLU GCNs. Xu et al. (2018) considered the case of ReLU nonlinearity, so that the update rule is $\bm{Z}^{\left(\ell\right)}=\mathsf{ReLU}(\hat{\bm{A}}\bm{Z}^{\left(\ell-1\right)}\bm{W}^{\left(\ell\right)})$. Additionally, it makes the simplifying assumption that each path in the computational graph is active with a fixed probability, $\rho$ (Kawaguchi, 2016, Assumption A1p-m). Accordingly, the sensitivity (in expectation) between any two nodes is given as
$\left\|\mathbb{E}_{\mathsf{ReLU}}\left[\frac{\partial\bm{z}^{\left(L\right)}_{i}}{\partial\bm{x}_{j}}\right]\right\|_{1}=\left[\rho\left\|\prod_{\ell=1}^{L}\bm{W}^{\left(\ell\right)}\right\|_{1}\right]\left(\hat{\bm{A}}^{L}\right)_{ij}=\zeta_{1}^{\left(L\right)}\left(\hat{\bm{A}}^{L}\right)_{ij}$ (3.8)
where $\zeta_{1}^{\left(L\right)}$ is independent of the choice of nodes $i,j\in\mathcal{V}$. Taking an expectation w.r.t. the random edge masks, we get
$\mathbb{E}_{\mathbf{M}^{\left(1\right)},\ldots,\mathbf{M}^{\left(L\right)}}\left[\left\|\mathbb{E}_{\mathsf{ReLU}}\left[\frac{\partial\bm{z}^{\left(L\right)}_{i}}{\partial\bm{x}_{j}}\right]\right\|_{1}\right]$ $=\zeta_{1}^{\left(L\right)}\mathbb{E}_{\mathbf{M}^{\left(1\right)},\ldots,\mathbf{M}^{\left(L\right)}}\left[\left(\prod_{\ell=1}^{L}\hat{\bm{A}}^{\left(\ell\right)}\right)_{ij}\right]$ (3.9)
$=\zeta_{1}^{\left(L\right)}\left(\mathbb{E}\left[\hat{\bm{A}}\right]^{L}\right)_{ij}$ (3.10)
Using Theorem 3.1, we conclude that in a ReLU-GCN, DropEdge, DropAgg and DropGNN will reduce the expected sensitivity between nodes L-hops away. Empirical observations in Figure 1 and Figure 4 suggest that we may expect an increase in sensitivity to neighboring nodes, but a significant decrease in sensitivity to those farther away.
Source-only Message Functions. Black et al. (2023, Lemma 3.2) considers MPNNs with aggregation functions of the form
$\mathsf{Agg}^{\left(\ell\right)}\left(\bm{z}_{i}^{\left(\ell-1\right)},\left\{\bm{z}_{j}^{\left(\ell-1\right)}:j\in\mathbb{S}^{\left(1\right)}\left(i\right)\right\}\right)=\sum_{j\in\mathbb{B}^{\left(1\right)}\left(i\right)}\hat{\bm{A}}_{ij}\mathsf{Msg}^{\left(\ell\right)}\left(\bm{z}_{j}^{\left(\ell-1\right)}\right)$ (3.11)
and $\mathsf{Upd}$ and $\mathsf{Msg}$ functions with bounded gradients. In this case, the sensitivity between two nodes $i,j\in\mathcal{V}$ can be bounded as
$\left\|\frac{\partial\bm{z}^{\left(L\right)}_{i}}{\partial\bm{x}_{j}}\right\|_{1}\leq\zeta_{2}^{\left(L\right)}\left(\sum_{\ell=0}^{L}\hat{\bm{A}}^{\ell}\right)_{ij}$ (3.12)
As before, we can use the independence of edge masks to get an upper bound on the expected sensitivity:
$\mathbb{E}_{\mathbf{M}^{\left(1\right)},\ldots,\mathbf{M}^{\left(L\right)}}\left[\left\|\frac{\partial\bm{z}^{\left(L\right)}_{i}}{\partial\bm{x}_{j}}\right\|_{1}\right]$ $\leq\zeta_{2}^{\left(L\right)}\left(\mathbb{E}_{\mathbf{M}^{\left(1\right)},\ldots,\mathbf{M}^{\left(L\right)}}\left[I_{N}+\sum_{\ell=1}^{L}\prod_{k=1}^{\ell}\hat{\bm{A}}^{\left(k\right)}\right]\right)_{ij}$ (3.13)
$=\zeta_{2}^{\left(L\right)}\left(\sum_{\ell=0}^{L}\mathbb{E}\left[\hat{\bm{A}}\right]^{\ell}\right)_{ij}$ (3.14)
Figure 5 shows the plot of the entries of $\sum_{\ell=0}^{6}\dot{\bm{P}}^{\ell}$ (i.e. for DropEdge), as in the upper bound above, with $\hat{\bm{A}}=\hat{\bm{A}}^{\text{asym}}$. We observe that the sensitivity between nearby nodes marginally increases, while that between distant nodes notably decreases (similar to Figure 1), suggesting significant over-squashing. Similar observations can be made with $\hat{\bm{A}}=\hat{\bm{A}}^{\text{sym}}$, and for other DropEdge-variants.
Source-and-Target Message Functions. Topping et al. (2022, Lemma 1) showed that if the aggregation function is instead given by
$\mathsf{Agg}^{\left(\ell\right)}\left(\bm{z}_{i}^{\left(\ell-1\right)},\left\{\bm{z}_{j}^{\left(\ell-1\right)}:j\in\mathbb{S}^{\left(1\right)}\left(i\right)\right\}\right)=\sum_{j\in\mathbb{B}^{\left(1\right)}\left(i\right)}\hat{\bm{A}}_{ij}\mathsf{Msg}^{\left(\ell\right)}\left(\bm{z}_{i}^{\left(\ell-1\right)},\bm{z}_{j}^{\left(\ell-1\right)}\right)$ (3.15)
then the sensitivity between nodes $i\in\mathcal{V}$ and $j\in\mathbb{S}^{\left(L\right)}\left(i\right)$ can be bounded as
$\left\|\frac{\partial\bm{z}^{\left(L\right)}_{i}}{\partial\bm{x}_{j}}\right\|_{1}\leq\zeta_{3}^{\left(L\right)}\left(\hat{\bm{A}}^{L}\right)_{ij}$ (3.16)
With random edge-dropping, this bound can be adapted as follows:
$\mathbb{E}_{\mathbf{M}^{\left(1\right)},\ldots,\mathbf{M}^{\left(L\right)}}\left[\left\|\frac{\partial\bm{z}^{\left(L\right)}_{i}}{\partial\bm{x}_{j}}\right\|_{1}\right]\leq\zeta_{3}^{\left(L\right)}\left(\mathbb{E}\left[\hat{\bm{A}}\right]^{L}\right)_{ij}$ (3.17)
which is similar to Equation 3.10, only with a different proportionality constant, that is anyway independent of the choice of nodes. Here, again, we invoke Theorem 3.1 to conclude that $(\mathbb{E}[\hat{\bm{A}}]^{L})_{ij}$ decreases monotonically with increasing DropEdge probability $q$. This implies that, in a non-linear MPNN with $\hat{\bm{A}}=\hat{\bm{A}}^{\text{asym}}$, DropEdge lowers the sensitivity bound given above. Empirical results in Figure 4 support the same conclusion for $\hat{\bm{A}}=\hat{\bm{A}}^{\text{sym}}$.
Message of the Section: Studying the expected propagation matrix allows us to predict the effect of random edge-dropping methods on information propagation in MPNNs. Specifically, DropEdge, DropAgg and DropGNN increase the sensitivity of nodes to their neighbors, but decrease it to nodes farther off. This suggests that these methods would be unsuitable for long-range tasks, where it is imperative to facilitate communication between distant nodes.
4 Experiments
Our theoretical analysis indicates that DropEdge may degrade the performance of GNNs on tasks that depend on capturing LRIs. In this section, we test this hypothesis by evaluating DropEdge models on both short-range and long-range tasks.
4.1 Setup
Datasets. Although identifying whether a task requires modeling LRIs can be challenging, understanding the structure of the datasets can provide some insight. For example, homophilic datasets have local consistency in node labels, i.e. nodes closely connected to each other have similar labels. On the other hand, in heterophilic datasets, nearby nodes often have dissimilar labels. Since DropEdge-variants increase the sensitivity of a node’s representations to its immediate neighbors, and reduce its sensitivity to distant nodes, we expect it to improve performance on homophilic datasets but harm performance on heterophilic ones (see Topping et al. (2022)). In this work, we use Cora (McCallum et al., 2000) and CiteSeer (Giles et al., 1998) as representatives of homophilic datasets (Zhu et al., 2020), and Squirrel, Chameleon and TwitchDE (Rozemberczki et al., 2021) to represent heterophilic datasets (Lim et al., 2021). The networks’ statistics are presented in Table 1, where we can note the significantly lower homophily measures of heterophilic datasets.
Dataset Nodes Edges Features Classes Homophily
Cora 2,708 10,556 1,433 7 0.766
CiteSeer 3,327 9,104 3,703 6 0.627
Chameleon 2,277 36,051 2,325 5 0.062
Squirrel 5,201 216,933 2,089 5 0.025
TwitchDE 9,498 306,276 128 2 0.142
Table 1: Dataset statistics. Number of edges excludes self-loops. Homophily measures from Lim et al. (2021).
Models. We use two MPNN architectures, GCN and GAT, to demonstrate the effect of DropEdge. GCN satisfies the model assumptions made in all the theoretical results presented in Section 3, while GAT does not fit into any of them. Therefore, GCN and GAT together provide a broad representation of different MPNN architectures. For GAT, we use 2 attentions heads in order to keep the computational load manageable, while at the same time harnessing the expressiveness of the multi-headed self-attention mechanism. As for the model depth, we vary it as $L=2,4,6,8$.
Dropout Methods. In addition to the DropEdge and its variants – DropNode, DropAgg and DropGNN – discussed in Section 3, we evaluate related stochastic regularization methods like regular dropout (Srivastava et al., 2014) (referred as Dropout), DropMessage (Fang et al., 2023) and SkipNode Lu et al. (2024). This allows us to study the pervasiveness of the problem of over-squashing over a large set of methods designed for alleviating over-smoothing.
Dropping Probability. For all the methods, the dropping probabilities are varied as $q=0.0,0.1,\ldots,0.9$, so as to reliably capture the trends in model accuracy. We adopt the common practice of turning the methods off at test-time ($q=0$), isolating their effect on optimization and generalization, which our theory does not address.
Other experimental details are reported in Appendix E. We conduct 5 independent runs for each dataset--model--dropout--probability configuration and report the average test accuracy (with early-stopping using the validation set).
4.2 Results
Dropout GNN Homophilic Heterophilic
Cora CiteSeer Chameleon Squirrel TwitchDE
DropEdge GCN $+0.396$ $+0.448$ $-0.723$ $-0.614$ $-0.509$
GAT $+0.547$ $+0.548$ $-0.236$ $-0.641$ $-0.390$
DropNode GCN $-0.320$ $-0.662$ $-0.863$ $-0.790$ $-0.465$
GAT $-0.706$ $-0.735$ $-0.855$ $-0.286$ $-0.512$
DropAgg GCN $-0.015$ $+0.294$ $-0.334$ $-0.317$ $-0.566$
GAT $-0.019$ $+0.392$ $-0.353$ $-0.398$ $-0.022$
DropGNN GCN $+0.360$ $+0.468$ $-0.746$ $-0.600$ $-0.671$
GAT $+0.236$ $+0.525$ $-0.507$ $-0.304$ $-0.139$
Dropout GCN $+0.034$ $-0.309$ $-0.610$ $-0.465$ $-0.300$
GAT $+0.120$ $-0.110$ $-0.836$ $-0.545$ $-0.397$
DropMessage GCN $-0.220$ $-0.607$ $-0.404$ $-0.364$ $-0.199$
GAT $+0.028$ $-0.368$ $-0.530$ $-0.697$ $-0.148$
SkipNode GCN $+0.161$ $+0.017$ $-0.528$ $-0.285$ $-0.244$
GAT $+0.403$ $+0.466$ $-0.436$ $+0.248$ $+0.278$
Table 2: Spearman correlation between dropping probability and test accuracy. Note the positive correlations for homophilic datasets and negative correlations for heterophilic datasets. Results suggest that random edge-dropping, although effective at improving generalization in short-range tasks, is unsuitable for long-range ones.
In Table 2, we report the rank correlation between the dropping probability and test accuracy of different dataset--model--dropout combinations. The statistics are computed over $5010\ (q=0.0,0.1,\ldots,0.9)\times 5\ (\text{samples})=$ 50 runs for each $L=2,4,6,8$, and then averaged. It is clear to see that in most combinations with the homophilic datasets Cora and CiteSeer, the correlation is positive, indicating that DropEdge and its variants improve test-time performance at short-range tasks. On the other hand, with the heterophilic datasets Chameleon, Squirrel and TwitchDE, the correlation values are negative. This suggests that the edge-dropping methods harm generalization in long-range tasks by forcing the model to overfit to short-range signals.
(a) Homophilic datasets
(b) Heterophilic datasets
Figure 2: Dropping probability versus test accuracy of DropEdge-GCN. The theory the explains the contrasting trends as follows: random edge-dropping pushes models to fit to local information during training, which is suitable for short-range tasks, but harms test-time performance in long-range ones.
Figure 2 allows us to visualize these trends for the homophilic and heterophilic datasets. For conciseness, we only display the results for DropEdge-GCNs. Clearly, on Cora and CiteSeer, the model’s test-time performance improves with increasing dropping probability, as has been reported earlier. However, on Chameleon, Squirrel and TwitchDE, the performance degrades with increasing dropping probability, as was suggested by our theoretical results. This highlights an important gap in our understanding of dropout methods – while their positive effects on model performance have been well-studied, making them popular choices for training deep GNNs, their evaluation has been limited to short-range tasks, leaving their negative impact on capturing LRIs overlooked.
Remark on DropNode. In Equation 3.4, we noted that DropNode does not suffer from loss in sensitivity. However, note that those results were in expectation. Moreover, our analysis did not account for the effects on the learning trajectory. In practice, a high DropNode probability would make it hard for information in the node features to reach distant nodes. This would prevent the model from learning to effectively combine information from large neighborhoods, harming generalization. In fact, in Table 2, we can see that it is the only dropping method with a negative correlation between dropping probability and test accuracy for each dataset--model combination, including homophilic ones. See Figure 6 for a visualization of its negative effects on test accuracy.
4.3 Over-squashing or Under-fitting?
Figure 3: DropEdge probability versus training accuracy of GCNs. The training performance improves with $q$, suggesting that the models are not underfitting. Instead, the reason for poor test-time performance (2(a)) is that models are over-fitting to short-range signals during training, resulting in poor generalization.
The results in the previous subsection suggest that using random edge-dropping to regularize model training leads to poor test-time performance. We hypothesize that this occurs because the models struggle to propagate information over long distances, causing overfitting of node representations to local neighborhoods. However, a confounding factor is at play: DropEdge variants reduce the generalization gap by preventing overfitting to the training set, which manifests as reduced training performance. If this regularization is too strong, it could lead to underfitting, which could also explain the poor test-time performance on heterophilic datasets. This concern is particularly relevant because the heterophilic networks are much larger than homophilic ones (Table 1), making them more prone to underfitting. To investigate this, we plot the training accuracies of deep DropEdge-GCNs on the heterophilic datasets; Figure 3 shows the results. It is clear that the models do not underfit as the dropping probability increases. In fact, somewhat unexpectedly, the training metrics improve. Together with the results in Figure 2, we conclude that DropEdge-like methods are detrimental in long-range tasks, as they cause overfitting to short-range artifacts in the training data, resulting in poor generalization at test-time.
Message of the Section: Random edge-dropping algorithms reduce the receptive field of MPNNs, forcing them to fit to short-range signals in the training set. While this may make deep GNNs suitable for homophilic datasets, it results in overfitting on heterophilic datasets, which leads to poor test-time performance. Therefore, these methods should be used only after carefully understanding the task at hand.
5 Conclusion
Our analysis points out a key assumption in algorithms designed for training deep GNNs: the idea that if a deep GNN is trainable, it must be able to model LRIs. Our results suggest that this, in fact, need not be true – we theoretically and empirically show that DropEdge-like algorithms exacerbate the over-squashing problem in deep GNNs, and degrade their performance on long-range tasks. Our results highlight a need for a thorough evaluation of methods employed when training deep GNNs, with regards to their capacity to capture LRIs. This will allow us to reliably deploy them in real-life, since we can be assured that the models did not simply overfit to short-range signals.
Limitations. While our theoretical analysis successfully predicts how DropEdge-variants affect test performance on short-range and long-range tasks, it is based on several simplifying assumptions. These assumptions, although standard in the literature, limit the generalizability of our conclusions to other architectures. Specifically, our analysis focuses on certain classes of MPNNs, excluding several GNN architectures specifically designed to enhance long-distance information propagation (see Singh (2024) for a review).
Practical Considerations. Previous studies have shown that random edge-dropping algorithms can effectively enhance generalization performance in short-range tasks, and our findings support this conclusion. However, we have also demonstrated that such algorithms can negatively impact over-squashing in MPNNs and harm test-time performance in long-range tasks. Therefore, we recommend exercising caution when using such methods, as careless application can result in models that generalize poorly, which can be detrimental in critical applications.
Future Directions. This work focuses on node-classification tasks, but it is also important to understand the effect of random edge-dropping in other practical settings, e.g. link prediction, and even graph-level tasks. In general, there is a need for a broader investigation into methods designed for training deep GNNs. Specifically, analyzing various strategies designed for mitigating over-smoothing (see Rusch et al. (2023) for a review), particularly in the context of over-squashing, could be invaluable for designing deep GNNs for long-range tasks.
Appendix A Related Works
A.1 Methods for Alleviating Over-smoothing
A popular choice for reducing over-smoothing in GNNs is to regularize the model. Recall that DropEdge (Rong et al., 2020) implicitly regularizes the model by adding noise to the learning trajectory (Section 2.2). Similarly, Dropout Srivastava et al. (2014) drops entries from the input features to each layer, and DropMessage (Fang et al., 2023) drops entries from the message matrix (before the aggregation step). Graph Drop Connect (GDC) (Hasanzadeh et al., 2020) combines DropEdge and DropMessage together, resulting in a layer-wise sampling scheme that uses a different subgraph for message-aggregation over each feature dimension. These methods successfully addressed the over-smoothing problem, enabling the training of deep GNNs, and performed competitively on several benchmarking datasets.
Another powerful form of implicit regularization is feature normalization, which has proven crucial in enhancing the performance and stability of several types of neural networks (Huang et al., 2020). Exploiting the inductive bias in graph-structured data, normalization techniques like PairNorm (Zhao & Akoglu, 2020), Differentiable Group Normalization (DGN) (Zhou et al., 2020) and NodeNorm (Zhou et al., 2021b) have been proposed to reduce over-smoothing in GNNs. On the other hand, Energetic Graph Neural Networks (EGNNs) Zhou et al. (2021a) explicitly regularize the optimization by constraining the layer-wise Dirichlet energy to a predefined range.
In a different vein, motivated by the success of residual networks (ResNets) (He et al., 2016) in computer vision, Li et al. (2019) proposed the use of residual connections to prevent the smoothing of representations. Residual connections successfully improved the performance of GCN on a range of graph-learning tasks. Chen et al. (2020) introduced GCNII, which uses skip connections from the input to all hidden layers. This layer wise propagation rule has allowed for training of ultra-deep networks – up to 64 layers. Some other architectures, like the Jumping Knowledge Network (JKNet) Xu et al. (2018) and the Deep Adaptive GNN (DAGNN) Liu et al. (2020), aggregate the representations from all layers, $\{\bm{z}^{\left(\ell\right)}_{i}\}_{\ell=1}^{L}$, before processing them through a readout layer.
A.2 Methods for Alleviating Over-squashing
In this section, we will review some of the graph rewiring methods proposed to address the problem of over-squashing. Particularly, we wish to emphasize a commonality among these methods – edge addition is necessary. As a reminder, graph rewiring refers to modifying the edge set of a graph by adding and/or removing edges in a systematic manner. In a special case, which includes many of the rewiring techniques we will discuss, the original topology is completely discarded, and only the rewired graph is used for message-passing.
Spatial rewiring methods use the topological relationships between the nodes in order to come up with a rewiring strategy. That is the graph rewiring is guided by the objective of optimizing some chosen topological properties. For instance, Alon & Yahav (2021) introduced a fully-adjacent (FA) layer, wherein messages are passed between all nodes. GNNs using a FA layer in the final message-passing step were shown to outperform the baselines on a variety of long-range tasks, revealing the importance of information exchange between far-off nodes which standard message-passing cannot facilitate. Topping et al. (2022) proposed a curvature-based rewiring strategy, called the Stochastic Discrete Ricci Flow (SDRF), which aims to reduce the “bottleneckedness” of a graph by adding suitable edges, while simultaneously removing edges in an effort to preserve the statistical properties of the original topology. Black et al. (2023) proposed the Greedy Total Resistance (GTR) technique, which optimizes the graph’s total resistance by greedily adding edges to achieve the greatest improvement. One concern with graph rewiring methods is that unmoderated densification of the graph, e.g. using a fully connected graph for propagating messages, can result in a loss of the inductive bias the topology provides, potentially leading to over-fitting. Accordingly, Gutteridge et al. (2023) propose a Dynamically Rewired (DRew) message-passing framework that gradually densifies the graph. Specifically, in a given layer $\ell$, node $i$ aggregates messages from its entire $\ell$-hop receptive field instead of just the immediate neighbors. This results in an improved communication over long distances while also retaining the inductive bias of the shortest distance between nodes.
[…]
Some other rewiring techniques don’t exactly classify as spatial or spectral methods. For instance, Probabilistically Rewired MPNN (PR-MPNN) Qian et al. (2024) learns to probabilistically rewire a graph, effectively mitigating under-reaching as well as over-squashing. Finally, Gabrielsson et al. (2023) proposed connecting all nodes at most $r$-hops away, for some $r\in\mathbb{N}$, and introducing positional embeddings to allow for distance-aware aggregation of messages.
A.3 Towards a Unified Treatment
Several studies have shown that an inevitable trade-off exists between the problems of over-smoothing and over-squashing, meaning that optimizing for one will compromise the other. For instance, Topping et al. (2022); Nguyen et al. (2023) showed that negatively curved edges create bottlenecks in the graph resulting in over-squashing of information. On the other hand, Nguyen et al. (2023, Proposition 4.3) showed that positively curved edges in a graph contribute towards the over-smoothing problem. To address this trade-off, they proposed Batch Ollivier-Ricci Flow (BORF), which adds new edges adjacent to the negatively curved ones, and simultaneously removes positively curved ones. In a similar vein, Giraldo et al. (2023) demonstrated that the minimum number of message-passing steps required to reach a given level of over-smoothing is inversely related to the Cheeger’s constant, $h_{\mathcal{G}}$. This again implies an inverse relationship between over-smoothing and over-squashing. To effectively alleviate the two issues together, they proposed the Stochastic Jost and Liu Curvature Rewiring (SJLR) algorithm, which adds edges that result in high improvement in the curvature of existing edges, while simultaneously removing those that have low curvature.
Despite the well-established trade-off between over-smoothing and over-squashing, some works have successfully tackled them together despite only adding or removing edges. One such work is Karhadkar et al. (2023), which proposed a rewiring algorithm that adds edges to the graph but does not remove any. The First-order Spectral Rewiring (FoSR) algorithm computes, as the name suggests, a first order approximation to the spectral gap of $\bm{L}^{\text{sym}}$, and adds edges with the aim of maximizing it. Since the spectral gap directly relates to Cheeger’s constant through Cheeger’s inequality (Alon & Milman, 1985; Alon, 1986; Sinclair & Jerrum, 1989), this directly decreases the over-squashing levels. Moreover, Karhadkar et al. (2023, Figure 5) empirically demonstrated that addition of (up to a small number of) edges selected by FoSR can lower the Dirichlet energy of the representations. Taking a completely opposite approach, CurvDrop (Liu et al., 2023) adapted DropEdge to remove negatively curved edges sampled from a distribution proportional to their curvatures. CurvDrop directly reduces over-squashing and, as a side benefit of operating on a sparser subgraph, also mitigates over-smoothing.
Appendix B Proofs
B.1 Expected Propagation Matrix under DropEdge
Lemma.
When using DropEdge, the expected propagation matrix is given as:
$\mathbb{E}_{\mathsf{DE}}\left[\hat{\bm{A}}^{\left(1\right)}_{ii}\right]$ $=\frac{1-q^{d_{i}+1}}{\left(1-q\right)\left(d_{i}+1\right)}$
$\mathbb{E}_{\mathsf{DE}}\left[\hat{\bm{A}}^{\left(1\right)}_{ij}\right]$ $=\frac{1}{d_{i}}\left(1-\frac{1-q^{d_{i}+1}}{\left(1-q\right)\left(d_{i}+1\right)}\right)$
where $\left(j\to i\right)\in\mathcal{E}$; $\dot{\bm{P}}_{ij}=0$ otherwise.
Proof.
Recall that under DropEdge, a self-loop is added to the graph after the edges are dropped, and then the normalization is performed. In other words, the self-loop is never dropped. Therefore, given the i.i.d. masks, $m_{1},\ldots,m_{d_{i}}\sim\text{Bern}\left(1-q\right)$, on incoming edges to node $i$, the total number of messages is given by
$1+\sum_{k=1}^{d_{i}}m_{k}=1+M_{i}$
where $M_{i}\sim\text{Binom}\left(d_{i},1-q\right)$. Under asymmetric normalization (see Section 2.1), the expected weight of the message along the self-loop is computed as follows:
$\mathbb{E}_{\mathsf{DE}}\left[\hat{\bm{A}}^{\left(1\right)}_{ii}\right]$ $=\mathbb{E}_{m_{1},\ldots,m_{d_{i}}}\left[\frac{1}{1+\sum_{k=1}^{d_{i}}m_{k}}\right]$ (B.1)
$=\mathbb{E}_{M_{i}}\left[\frac{1}{1+M_{i}}\right]$ (B.2)
$=\sum_{k=0}^{d_{i}}\binom{d_{i}}{k}\left(1-q\right)^{k}\left(q\right)^{d_{i}-k}\left(\frac{1}{1+k}\right)$ (B.3)
$=\frac{1}{\left(1-q\right)\left(d_{i}+1\right)}\sum_{k=0}^{d_{i}}\binom{d_{i}+1}{k+1}\left(1-q\right)^{k+1}\left(q\right)^{d_{i}-k}$ (B.4)
$=\frac{1}{\left(1-q\right)\left(d_{i}+1\right)}\sum_{k=1}^{d_{i}+1}\binom{d_{i}+1}{k}\left(1-q\right)^{k}\left(q\right)^{d_{i}+1-k}$ (B.5)
$=\frac{1-q^{d_{i}+1}}{\left(1-q\right)\left(d_{i}+1\right)}$ (B.6)
Similarly, if the Bernoulli mask corresponding to $j\to i$ is $1$, then the total number of incoming messages to node $i$ is given by
$2+\sum_{k=1}^{d_{i}-1}m_{k}$
including one self-loop, which is never dropped, as noted earlier. On the other hand, the weight of the edge is simply 00 if the corresponding Bernoulli mask is 00. Using the Law of Total Expectation, the expected weight of the edge $j\to i$ can be computed as follows:
$\mathbb{E}_{\mathsf{DE}}\left[\hat{\bm{A}}^{\left(1\right)}_{ij}\right]$ $=q\cdot 0+\left(1-q\right)\mathbb{E}_{m_{1},\ldots,m_{d_{i}-1}}\left[\frac{1}{2+\sum_{k=1}^{d_{i}-1}m_{k}}\right]$ (B.7)
$=\left(1-q\right)\sum_{k=0}^{d_{i}-1}\binom{d_{i}-1}{k}\left(1-q\right)^{k}\left(q\right)^{d_{i}-1-k}\left(\frac{1}{2+k}\right)$ (B.8)
$=\sum_{k=0}^{d_{i}-1}\frac{\left(d_{i}-1\right)!}{\left(k+2\right)!\left(d_{i}-1-k\right)!}\left(1-q\right)^{k+1}\left(q\right)^{d_{i}-1-k}\left(k+1\right)$ (B.9)
$=\sum_{k=2}^{d_{i}+1}\frac{\left(d_{i}-1\right)!}{\left(k\right)!\left(d_{i}+1-k\right)!}\left(1-q\right)^{k-1}\left(q\right)^{d_{i}+1-k}\left(k-1\right)$ (B.10)
$=\frac{1}{d_{i}\left(d_{i}+1\right)\left(1-q\right)}\sum_{k=2}^{d_{i}+1}\binom{d_{i}+1}{k}\left(1-q\right)^{k}\left(q\right)^{d_{i}+1-k}\left(k-1\right)$ (B.11)
$=\frac{1}{d_{i}\left(d_{i}+1\right)\left(1-q\right)}\left[\left(d_{i}+1\right)\left(1-q\right)-1+q^{d_{i}+1}\right]$ (B.12)
$=\frac{1}{d_{i}}\left(1-\mathbb{E}_{\mathsf{DE}}\left[\hat{\bm{A}}^{\left(1\right)}_{ii}\right]\right)$ (B.13)
B.2 Sensitivity in L-layer Linear GCNs
Theorem.
In an L-layer linear GCN with $\hat{\bm{A}}=\hat{\bm{A}}^{\text{asym}}$, using DropEdge, DropAgg or DropGNN decreases the sensitivity of a node $i\in\mathcal{V}$ to another node $j\in\mathbb{S}^{\left(L\right)}\left(i\right)$, thereby reducing its effective receptive field.
$\mathbb{E}_{\ldots}\left[\left(\hat{\bm{A}}^{L}\right)_{ij}\right]=\sum_{\left(u_{0},\ldots,u_{L}\right)\in\mathsf{Paths}\left(j\to i\right)}\prod_{\ell=1}^{L}\mathbb{E}_{\ldots}\left[\hat{\bm{A}}_{u_{\ell}u_{\ell-1}}\right]<\mathbb{E}_{\mathsf{ND}}\left[\left(\hat{\bm{A}}^{L}\right)_{ij}\right]$ (B.14)
where $\mathsf{ND}$ refers to a NoDrop model ($q=0$), the placeholder $\cdots$ can be replaced with one of the edge-dropping methods $\mathsf{DE}$, $\mathsf{DA}$ or $\mathsf{DG}$, and the corresponding entries of $\mathbb{E}_{\ldots}[\hat{\bm{A}}]$ can be plugged in from Equation 3.3, Equation 3.5 and Equation 3.6, respectively. Moreover, the sensitivity monotonically decreases as the dropping probability is increased.
Proof.
Recall that $\dot{\bm{P}}$ can be viewed as the transition matrix of a non-uniform random walk, such that $\dot{\bm{P}}_{uv}=\mathbb{P}\left(u\to v\right)$. Intuitively, since there is no self-loop on any given L-length path connecting nodes $i$ and $j$ (which are assumed to be L-hops away), the probability of each transition on any path connecting these nodes is reduced. Therefore, so is the total probability of transitioning from $i$ to $j$ in exactly L hops.
More formally, denote the set of paths connecting the two nodes by
$\mathsf{Paths}\left(j\to i\right)=\left\{\left(u_{0},\ldots,u_{L}\right):u_{0}=j;u_{L}=i;\left(u_{\ell-1}\to u_{\ell}\right)\in\mathcal{E},\forall\ell\in\left[L\right]\right\}$
The $\left(i,j\right)$-entry in the propagation matrix is given by
$\left(\dot{\bm{P}}^{L}\right)_{ij}$ $=\sum_{\left(u_{0},\ldots,u_{L}\right)\in\mathsf{Paths}\left(j\to i\right)}\prod_{\ell=1}^{L}\dot{\bm{P}}_{u_{\ell}u_{\ell-1}}$ (B.15)
Since there is no self-loop on any of these paths,
$\left(\dot{\bm{P}}^{L}\right)_{ij}$ $=\sum_{\left(u_{0},\ldots,u_{L}\right)\in\mathsf{Paths}\left(j\to i\right)}\prod_{\ell=1}^{L}\frac{1}{d_{u_{\ell}}}\left(1-\frac{1-q^{d_{u_{\ell}}+1}}{\left(1-q\right)\left(d_{u_{\ell}}+1\right)}\right)$ (B.16)
$<\sum_{\left(u_{0},\ldots,u_{L}\right)\in\mathsf{Paths}\left(j\to i\right)}\prod_{\ell=1}^{L}\left(\frac{1}{d_{u_{\ell}}+1}\right)$ (B.17)
The right hand side of the inequality is the $\left(i,j\right)$-entry in the Lth power of the propagation matrix of a NoDrop model. From Equation 3.5 and Equation 3.6, we know that Equation B.17 is true for DropAgg and DropGNN as well. We conclude the first part of the proof using Equation 3.2 – the sensitivity of node $i$ to node $j$ is proportional to $(\dot{\bm{P}}^{L})_{ij}$.
Next, we recall the geometric series for any $q$:
$1+q+\ldots+q^{d}=\frac{1-q^{d+1}}{1-q}$ (B.18)
Each of the terms on the right are increasing in $q$, hence, all the $\dot{\bm{P}}_{u_{\ell}u_{\ell-1}}$ factors are decreasing in $q$. Using this result with Equation B.15, we conclude the second part of the theorem.
Appendix C Test-time Monte-Carlo Averaging
In Section 3, we focused on the expected sensitivity of the stochastic representations in models using DropEdge-like regularization strategies. This corresponds to their training-time behavior, wherein the activations are random. At test-time, the standard practice is to turn these methods off by setting $q=0$. However, this raises the over-smoothing levels back up Xuanyuan et al. (2023). Another way of making predictions is to perform multiple stochastic forward passes, as during training, and then averaging the model outputs. This is similar to Monte-Carlo Dropout, which is an efficient way of ensemble averaging in MLPs Gal & Ghahramani (2016b), CNNs Gal & Ghahramani (2016a) and RNNs Gal & Ghahramani (2016c). In addition to alleviating over-smoothing, this approach also outperforms the standard implementation Xuanyuan et al. (2023). We can study the effect of random edge-dropping in this setting by examining the sensitivity of the expected representations:
$\left\|\frac{\partial}{\partial\bm{x}_{j}}\mathbb{E}\left[\bm{z}^{\left(L\right)}_{i}\right]\right\|_{1}$
In linear models, the order of the two operations – expectation and sensitivity computation – is irrelevant:
$\mathbb{E}\left[\left\|\frac{\partial\bm{z}^{\left(L\right)}_{i}}{\partial\bm{x}_{j}}\right\|_{1}\right]=\left\|\mathbb{E}\left[\hat{\bm{A}}_{ij}\right]\bm{W}\right\|_{1}=\left\|\mathbb{E}\left[\frac{\partial\bm{z}^{\left(L\right)}_{i}}{\partial\bm{x}_{j}}\right]\right\|_{1}=\left\|\frac{\partial}{\partial\bm{x}_{j}}\mathbb{E}\left[\bm{z}^{\left(L\right)}_{i}\right]\right\|_{1}$ (C.1)
In general, the two quantities can be related using the convexity of norms and Jensen’s inequality:
$\left\|\frac{\partial}{\partial\bm{x}_{j}}\mathbb{E}\left[\bm{z}^{\left(L\right)}_{i}\right]\right\|_{1}\leq\mathbb{E}\left[\left\|\frac{\partial\bm{z}^{\left(L\right)}_{i}}{\partial\bm{x}_{j}}\right\|_{1}\right]$ (C.2)
Therefore, the discussion in the previous subsections extends to the MC-averaged representations as well. Although tighter bounds may be derived for this setting, we will leave that for future works.
Appendix D Additional Figures
In this section, we present some additional figures that demonstrate the negative effects of random edge-dropping, particularly focusing on providing empirical evidence for scenarios not covered by the theory in Section 3. Additionally, we provide visual evidence of the negative effects of DropNode, despite the fact that it preserves sensitivity between nodes (in expectation).
D.1 Symmetrically Normalized Propagation Matrix
Figure 4: Entries of $\ddot{\bm{P}}^{6}$, averaged after binning node-pairs by their shortest distance.
The results in Section 3.1 correspond to the use of $\hat{\bm{A}}=\hat{\bm{A}}^{\text{asym}}$ for aggregating messages – in each message passing step, only the in-degree of node $i$ is used to compute the aggregation weights of the incoming messages. In practice, however, it is more common to use the symmetrically normalized propagation matrix, $\hat{\bm{A}}=\hat{\bm{A}}^{\text{sym}}$, which ensures that nodes with high degree do not dominate the information flow in the graph Kipf & Welling (2017). As in Equation 3.2, we are looking for
$\mathbb{E}_{\mathbf{M}^{\left(1\right)},\ldots,\mathbf{M}^{\left(L\right)}}\left[\prod_{l=1}^{L}\hat{\bm{A}}^{\left(\ell\right)}\right]=\ddot{\bm{P}}^{L}$
where $\ddot{\bm{P}}\coloneqq\mathbb{E}_{\mathsf{DE}}[\hat{\bm{A}}^{\text{sym}}]$. While $\ddot{\bm{P}}$ is analytically intractable, we can approximate it using Monte-Carlo sampling. Accordingly, we use 20 samples of $\mathbf{M}$ to compute an approximation of $\ddot{\bm{P}}$, and plot out the entries of $\ddot{\bm{P}}^{L}$, as we did for $\dot{\bm{P}}^{L}$ in Figure 1. The results are presented in Figure 4, which shows that while the sensitivity between nearby nodes is affected to a lesser extent compared to those observed in Figure 1, that between far-off nodes is significantly reduced, same as earlier.
D.2 Upper Bound on Expected Sensitivity
Figure 5: Entries of $\sum_{l=0}^{6}\dot{\bm{P}}^{l}$, averaged after binning node-pairs by their shortest distance.
Black et al. (2023) showed that the sensitivity between any two nodes in a graph can be bounded using the sum of the powers of the propagation matrix. In Section 3.2, we extended this bound to random edge-dropping methods with independent edge masks smapled in each layer:
$\mathbb{E}_{\mathbf{M}^{\left(1\right)},\ldots,\mathbf{M}^{\left(L\right)}}\left[\left\|\frac{\partial\bm{z}^{\left(L\right)}_{i}}{\partial\bm{x}_{j}}\right\|_{1}\right]=\zeta_{3}^{\left(L\right)}\left(\sum_{\ell=0}^{L}\mathbb{E}\left[\hat{\bm{A}}\right]^{\ell}\right)_{ij}$
Although this bound does not have a closed form, we can use real-world graphs to study its entries. We randomly sample 100 molecular graphs from the Proteins dataset Dobson & Doig (2003) and plot the entries of $\sum_{l=0}^{6}\dot{\bm{P}}^{\ell}$ (corresponding to DropEdge) against the shortest distance between node-pairs. The results are presented in Figure 5. We observe a polynomial decline in sensitivity as the DropEdge probability increases, suggesting that it is unsuitable for capturing LRIs.
D.3 Test Accuracy versus DropNode Probability
(a) Homophilic datasets
(b) Heterophilic datasets
Figure 6: Dropping probability versus test accuracy of DropNode-GCN.
In Equation 3.4, we noted that the expectation of sensitivity remains unchanged when using DropNode. However, these results were only in expectation. In practice, a high DropNode probability will result in poor communication between distant nodes, preventing the model from learning to effectively model LRIs. This is supported by the results in Table 2, where we observed a negative correlation between the test accuracy and DropNode probability. Moreover, DropNode was the only algorithm which recorded negative correlations on homophilic datasets. In Figure 6, we visualize these relationships, noting the stark contrast with Figure 2, particularly in the trends with homophilic datasets.
Appendix E Experiment Details
In this section, we expand on the details of the experiment in Section 4.
Descriptions of the Datasets. Cora McCallum et al. (2000) and CiteSeer Giles et al. (1998) are citation networks – their nodes represent scientific publications and an edge between two nodes indicates that one of them has cited the other. The features of each publication are represented by a binary vector, where each index indicates whether a specific word from a dictionary is present or absent. Several studies have showed that these datasets have high homophily in node labels Lim et al. (2021); Zhu et al. (2020) and that they are modelled much better by shallower networks than by deeper ones Zhou et al. (2020). Chameleon and Squirrel Rozemberczki et al. (2021) are networks of English Wikipedia web pages on the respective topics, and the edges between web pages indicate links between them. The task is to predict the average-monthly traffic on each of the web pages. Finally, TwitchDE Rozemberczki et al. (2021) is a network of Twitch users in Germany, with the edges between them representing their mutual follower relationships. The node features are embeddings of the games played by the users. The task is to predict whether the users use explicit language.
Training Hyperparameters. We standardise most of the hyperparameters across all experiments in order to isolate the effect of dropping probability. Specifically, we fix the size of the hidden representations in each layer at 64, and a linear readout layer is used to compute the node-level logits. The models are trained using the Adam optimizer Kingma & Ba (2015), with a learning rate of $3\times 10^{-3}$ and a weight decay of $5\times 10^{-4}$, for a total of 300 epochs. For GCN, we use symmetric normalization of the adjacency matrix to compute the edge weights Kipf & Welling (2017).
GAT Runs. It is quite problematic to train deep GAT models due to vanishing gradients (Dasoulas et al., 2021). Accordingly, we discard the runs where the model fails to learn, and performs just as well as a random classifier. Specifically, we compute the class distribution in each of the networks, and discard the runs where the test performance does not exceed the maximum proportion. This comes out to be $0.3021$, $0.2107$, $0.2288$, $0.2003$ and $0.6045$ for Cora, CiteSeer, Chameleon, Squirrel and TwitchDE, respectively.
E.1 Extension to GraphSAGE and ResGCN
Our exploration in this work has been mainly limited to stochastic regularization methods, including edge-dropping methods, and related ones like Dropout Srivastava et al. (2014), DropMessage Fang et al. (2023) and SkipNode Lu et al. (2024). We may follow a similar intuition to study other methods as well. For example, GraphSAGE Hamilton et al. (2017) operates by sampling a subset of neighbors for message aggregation in each step. From an over-squashing perspective, it is very similar to DropEdge and therefore, we can expect a similar decline in sensitivity between distant nodes when using GraphSAGE.
Figure 7: Influence distributions of a regular GCN and a ResGCN. Residual connections reduce the influence of nodes farther away by amplifying the influence of nodes close by.
On the other hand, residual connection reduce the influence of distant nodes by increasing the influence of nearby neighbors. For example, in the case of ResGCN, the operation $\bm{Z}^{\left(\ell+1\right)}=\hat{\bm{A}}\bm{Z}^{\left(\ell\right)}\bm{W}^{\left(\ell+1\right)}+\bm{Z}^{\left(\ell\right)}$ first collects information from nodes $(\ell+1)$-hops away, like in a regular GCN model, and then adds information from nodes up to $\ell$-hops away, effectively decreasing the influence of the nodes $(\ell+1)$-hops away. In other words, we may expect residual connections to be similarly unsuited for long-range tasks, not by reducing the strength of the signal over long distances, but rather by amplifying the strength of short-range signals. In Figure 7, we present the empirical influence distribution Xu et al. (2018) of a 6-layer ResGCN, computed using the Cora network. Here, we can clearly see that residual connections significantly reduce the influence of distant nodes on final representations, and can thus be expected to be unsuitable for long-range tasks.

### v2

@@
-MsgABD: EveryShipout initializing macros
+Jasraj Singh Independent Researcher &Keyue Jiang
 Abstract
-Message Passing Neural Networks (MPNNs) are a class of Graph Neural Networks (GNNs) that leverage the graph topology to propagate messages across increasingly larger neighborhoods.
-The message-passing scheme leads to two distinct challenges: over-smoothing and over-squashing.
-While several algorithms, e.g.
-DropEdge and its variants – DropNode, DropAgg and DropGNN – have successfully addressed the over-smoothing problem, their impact on over-squashing remains largely unexplored.
-This represents a critical gap in the literature as failure to mitigate over-squashing would make these methods unsuitable for long-range tasks.
-In this work, we take the first step towards closing this gap by studying the aforementioned algorithms in the context of over-squashing.
-We present novel theoretical results that characterize the negative effects of DropEdge on sensitivity between distant nodes, suggesting its unsuitability for long-range tasks.
-Our findings are easily extended to its variants, allowing us to build a comprehensive understanding of how they affect over-squashing.
-We evaluate these methods using real-world datasets, demonstrating their detrimental effects.
-Specifically, we show that while DropEdge-variants improve test-time performance in short-range tasks, they deteriorate performance in long-range ones.
-Our theory explains these results as follows: random edge-dropping lowers the effective receptive field of GNNs, which although beneficial for short-range tasks, misaligns the models on long-range ones.
-This forces the models to overfit to short-range artefacts in the training set, resulting in poor generalization.
-Our conclusions highlight the need to re-evaluate various methods designed for training deep GNNs, with a renewed focus on modelling long-range interactions.
+Message Passing Neural Networks (MPNNs) are a class of Graph Neural Networks (GNNs) that propagate information across the graph via local neighborhoods.
+The scheme gives rise to two key challenges: over-smoothing and over-squashing.
+While several Dropout-style algorithms, such as DropEdge and DropMessage, have successfully addressed over-smoothing, their impact on over-squashing remains largely unexplored.
+This represents a critical gap in the literature, as failure to mitigate over-squashing would make these methods unsuitable for long-range tasks – the intended use case of deep MPNNs.
+In this work, we study the aforementioned algorithms, and closely related edge-dropping algorithms – DropNode, DropAgg and DropGNN – in the context of over-squashing.
+We present theoretical results showing that DropEdge-variants reduce sensitivity between distant nodes, limiting their suitability for long-range tasks.
+To address this, we introduce DropSens, a sensitivity-aware variant of DropEdge that explicitly controls the proportion of information lost due to edge-dropping, thereby increasing sensitivity to distant nodes despite dropping the same number of edges.
+Our experiments on long-range synthetic and real-world datasets confirm the predicted limitations of existing edge-dropping and feature-dropping methods.
+Moreover, DropSens consistently outperforms graph rewiring techniques designed to mitigate over-squashing, suggesting that simple, targeted modifications can substantially improve a model’s ability to capture long-range interactions.
+Our conclusions highlight the need to re-evaluate and re-design existing methods for training deep GNNs, with a renewed focus on modelling long-range interactions.
 1 Introduction
-Graph-structured data is ubiquitous – it is found in social media platforms, online retail platforms, molecular structures, transportation networks, and even computer systems.
-Graph neural networks (GNNs) (Scarselli et al., 2009; Li et al., 2016) are powerful neural models developed for modelling graph-structured data, and have found applications in several real-world scenarios (Gao et al., 2018; You et al., 2020c; a; 2022; b; Ying et al., 2018; Zheng et al., 2022; Monti et al., 2017; Zitnik & Leskovec, 2017; Wale & Karypis, 2006).
-A popular class of GNNs, called message-passing neural networks (MPNNs) (Gilmer et al., 2017), recursively process neighborhood information using message-passing layers.
-These layers are stacked to allow each node to aggregate information from increasingly larger neighborhoods, akin to how convolutional neural networks (CNNs) learn hierarchical features for images (LeCun et al., 1989).
-However, unlike in image-based deep learning, where ultra-deep CNN architectures have led to performance breakthroughs (Szegedy et al., 2015; He et al., 2016), shallow GNNs often outperform deeper models on many graph learning tasks (Zhou et al., 2021b).
-This is because deep GNNs suffer from unique issues like over-smoothing (Oono & Suzuki, 2020) and over-squashing (Alon & Yahav, 2021), which makes training them notoriously difficult.
+Graph neural networks (GNNs) [71, 51] are powerful neural models developed for modelling graph-structured data, and have found applications in several real-world scenarios [29, 91, 89, 92, 90, 88, 94, 60, 99, 82].
+A popular class of GNNs, called message-passing neural networks (MPNNs) [32], recursively process neighborhood information using message-passing layers.
+These layers are stacked to allow each node to aggregate information from increasingly larger neighborhoods, akin to how convolutional neural networks (CNNs) learn hierarchical features for images [48].
+However, unlike in image-based deep learning, where ultra-deep CNN architectures have led to performance breakthroughs [78, 38], shallow GNNs often outperform deeper models on many graph learning tasks [97].
+This is because deep GNNs suffer from unique issues like over-smoothing [64] and over-squashing [4], which makes training them notoriously difficult.
 Over-smoothing refers to the problem of node representations becoming too similar as they are recursively processed.
 This is undesirable since it limits the GNN from effectively utilizing the information in the input features.
-The problem has garnered significant attention from the research community, resulting in a suite of algorithms designed to address it Rusch et al.
-(2023).
-Amongst these methods are a collection of random edge-dropping algorithms, including DropEdge (Rong et al., 2020) and its variants – DropNode (Feng et al., 2020), DropAgg Jiang et al.
-(2023) and DropGNN (Papp et al., 2021) – which act as message-passing reducers.
+The problem has garnered significant attention from the research community, resulting in a suite of algorithms designed to address it [70] (see Section A.1 for an overview of representative methods).
+Amongst these methods are a collection of random edge-dropping algorithms, including DropEdge [68], DropNode [23], DropAgg [43] and DropGNN [65] – which we will collectively refer to as DropEdge-variants – which act as message-passing reducers.
+In addition, we have DropMessage [21], which performs Dropout [77] on the message matrices, instead of the feature matrices; we will collectively refer to these two methods as Dropout-variants since they are applied along the feature dimensions.
 The other issue specific to GNNs is over-squashing.
-In certain graph structures, neighborhood size grows exponentially with distance from the source (Chen et al., 2018b), causing information to be lost as it passes through graph bottlenecks (Alon & Yahav, 2021).
+In certain graph structures, neighborhood size grows exponentially with distance from the source [12], causing information to be lost as it passes through graph bottlenecks [4].
 This limits MPNNs’ ability to enable communication between distant nodes, which is crucial for good performance on long-range tasks.
-To alleviate over-squashing, several graph-rewiring techniques have been proposed, which aim to improve graph connectivity by adding edges in a strategic manner [^1] Sometimes, along with removal of some edges to preserve statistical properties of the original topology.
-(Karhadkar et al., 2023; Deac et al., 2022; Black et al., 2023; Nguyen et al., 2023; Alon & Yahav, 2021).
+To alleviate over-squashing, several graph-rewiring techniques have been proposed, which aim to improve graph connectivity by adding edges in a strategic manner [44, 16, 8, 63, 4] (see Section A.3 for an overview of representative methods).
+[^1] Sometimes, along with removal of some edges to preserve statistical properties of the original topology.
 In contrast, the DropEdge-variants only remove edges, which should, in principle, amplify over-squashing levels.
-The empirical evidence in support of methods designed for training deep GNNs has been majorly collected on short-range tasks.
+The same can be intuitively argued about Dropout-variants.
+Empirical evidence in support of methods designed for training deep GNNs has been majorly collected on short-range tasks (see Section A.2 for a detailed discussion).
 That is, it simply suggests that these methods prevent loss of local information, but it remains inconclusive if they facilitate capturing long-range interactions (LRIs).
-Of course, on long-range tasks, deeper GNNs are useless if they cannot capture LRIs.
-This is especially a concern for DropEdge-variants since evidence suggests that alleviating over-smoothing with graph rewiring could exacerbate over-squashing (Nguyen et al., 2023; Giraldo et al., 2023).
-Theoretical Contributions.
-In this work, we uncover the effects of random edge-dropping algorithms on over-squashing in MPNNs.
-By explicitly computing the expected sensitivity of the node representations to the node features (Topping et al., 2022) (inversely related to over-squashing) in a linear GCN Kipf & Welling (2017), we show that these methods provably reduce the effective receptive field of the model.
-Precisely speaking, the rate at which sensitivity between distant nodes decays is polynomial w.r.t.
-the dropping probability.
-Finally, we extend the existing theoretical results on sensitivity in nonlinear MPNNs Black et al.
-(2023); Di Giovanni et al.
-(2023); Xu et al.
-(2018) to the random edge-dropping setting, again showing that these algorithms exacerbate the over-squashing problem.
-Experimental Results.
-We evaluate the DropEdge-variants on node classification tasks using GCN and GAT Veličković et al.
-(2018) architectures.
-Specifically, we assess their performance on homophilic datasets – Cora McCallum et al.
-(2000) and CiteSeer Giles et al.
-(1998) – which represent short-range tasks, and heterophilic datasets – Chameleon, Squirrel, TwitchDE Rozemberczki et al.
-(2021) – which correspond to long-range tasks.
-Our results indicate an increasing trend in test accuracy for homophilic datasets and a declining trend for heterophilic datasets as the dropping probability increases.
-Accordingly, we hypothesize that edge-dropping algorithms improve performance on short-range tasks, as has been reported earlier, by reducing the receptive field of the GNN and increasing model-dataset alignment, However, for long-range tasks, they decrease the model-dataset alignment, resulting in poor generalization.
+Of course, on long-range tasks, deep GNNs are useless if they cannot capture LRIs.
+This is especially a concern for DropEdge-variants since evidence suggests that alleviating over-smoothing with graph rewiring could exacerbate over-squashing [63, 34].
+Contributions.
+In this work, we precisely characterize the effects of random edge-dropping algorithms on over-squashing in MPNNs.
+By explicitly computing the expected sensitivity of the node representations to the node features [79] (inversely related to over-squashing) in a linear Graph Convolutional Network (GCN) [47], we show that these methods provably reduce the effective receptive field of the model.
+Precisely speaking, the rate at which sensitivity between nodes decays is exponential w.r.t.
+the distance between them.
+We also extend the existing theoretical results on sensitivity in nonlinear MPNNs [8, 18, 84] to the random edge-dropping setting, concluding that these algorithms exacerbate the over-squashing problem.
+We use our analysis of GCNs to design a sensitivity-aware DropEdge-variant, named DropSens, that enjoys the representational expressivity of DropEdge without suffering from over-squashing, thereby demonstrating how algorithms can be readily adapted for long-range tasks.
+We evaluated the DropEdge- and Dropout-variants on long-range datasets using GCN, Graph Isomorphism Network (GIN) [85] and Graph Attention Network (GAT) [81] architectures.
+Specifically, we follow the setup in Giovanni et al.
+[33] with the SyntheticZINC dataset, in Topping et al.
+[79] with real-world homophilic (corresponding to short-range tasks) and heterophilic (long-range tasks) node classification datasets, and in Black et al.
+[8], Karhadkar et al.
+[44] with graph classification datasets.
+Our results indicate that while the random dropping methods improve model performance in short-range tasks, they are often ineffective, and sometimes even detrimental, to long-range task performance.
+Finally, we present results for DropSens, which outperforms state-of-the-art graph rewiring methods aimed at addressing over-squashing at node classification and graph-classification tasks.
+These findings point to the importance of re-evaluating the methods used to train deep GNNs, especially in terms of how well they capture LRIs.
 2 Background
@@
 2.1 Graph Neural Networks
-Graph neural networks (GNNs) operate on inputs of the form $\left(\mathcal{G},\bm{X}\right)$, where $\mathcal{G}$ encodes the graph topology and $\bm{X}\in^{N\times H^{\left(0\right)}}$ collects the node features [^2] To keep things simple, we will ignore edge features..
-Message-passing neural networks (MPNNs) Gilmer et al.
-(2017) are a special class of GNNs which recursively aggregate information from the 1-hop neighborhood of each node using message-passing layers.
+Graph Neural Networks (GNNs) operate on inputs of the form $\left(\mathcal{G},\bm{X}\right)$, where $\mathcal{G}$ encodes the graph topology and $\bm{X}\in{}^{N\times H^{\left(0\right)}}$ collects the node features.
+[^2] To keep things simple, we will ignore edge features.
+Message-Passing Neural Networks (MPNNs) [32] are a special class of GNNs which recursively aggregate information from the 1-hop neighborhood of each node using message-passing layers.
 An L-layer MPNN is given as
@@
 where $\bm{Z}^{\left(0\right)}=\bm{X}$, $\mathsf{Agg}^{\left(\ell\right)}$ denotes the aggregation functions, $\mathsf{Upd}^{\left(\ell\right)}$ the update functions, and $\mathsf{Out}$ the readout function.
-Since the final representation of node $i$ is a function of the input features of nodes at most L-hops away from it, its receptive field is given by $\mathbb{B}^{\left(L\right)}\left(i\right)\coloneqq\left\{j\in\mathcal{V}:d_{\mathcal{G}}\left(j,i\right)\leq L\right\}$.
-For example, a GCN Kipf & Welling (2017) updates node representations as the weighted sum of its neighbors’ representations:
+Since $\bm{z}_{i}^{\left(L\right)}$ is a function of the input features of nodes at most L-hops away from it, its receptive field is given by $\mathbb{B}^{\left(L\right)}\left(i\right)\coloneqq\left\{j\in\mathcal{V}:d_{\mathcal{G}}\left(j,i\right)\leq L\right\}$.
+For example, a GCN [47] updates node representations as the weighted sum of its neighbors’ representations:
 $\bm{Z}^{\left(\ell\right)}=\sigma\left(\hat{\bm{A}}\bm{Z}^{\left(\ell-1\right)}\bm{W}^{\left(\ell\right)}\right)$ (2.2)
@@
 $\hat{\bm{A}}_{ij}\neq 0$ if and only if $\left(j\to i\right)\in\mathcal{E}$ or $i=j$, and $\bm{W}^{\left(\ell\right)}\in{}^{H^{\left(\ell-1\right)}\times H^{\left(\ell\right)}}$ is a weight matrix.
-The original choice for $\hat{\bm{A}}$ was the symmetrically normalized adjacency matrix $\hat{\bm{A}}^{\text{sym}}\coloneqq\widetilde{\bm{D}}^{-1/2}\widetilde{\bm{A}}\widetilde{\bm{D}}^{-1/2}$ Kipf & Welling (2017), where $\widetilde{\bm{A}}=\bm{A}+\bm{I}_{N}$ and $\widetilde{\bm{D}}=\text{diag}(\widetilde{\bm{A}}\mathbf{1}_{N})$.
-However, several influential works have also used the asymmetrically normalized adjacency, $\hat{\bm{A}}^{\text{asym}}\coloneqq\widetilde{\bm{D}}^{-1}\widetilde{\bm{A}}$ Hamilton et al.
-(2017); Schlichtkrull et al.
-(2017); Li et al.
-(2018).
-2.2 DropEdge
-DropEdge is a random data augmentation technique that works by sampling a subgraph of the original input graph in each layer, and uses that for message passing Rong et al. (2020):
-$\begin{split}\mathbf{M}^{\left(\ell\right)}&\sim\left\{\text{Bern}\left(1-q\right)\right\}^{N\times N}\\ \widetilde{\bm{A}}^{\left(\ell\right)}&=\mathbf{M}^{\left(\ell\right)}\circ\bm{A}+\bm{I}_{N}\end{split}$ (2.3)
+The original choice for $\hat{\bm{A}}$ was the symmetrically normalized adjacency matrix $\hat{\bm{A}}^{\text{sym}}\coloneqq\widetilde{\bm{D}}^{-1/2}\widetilde{\bm{A}}\widetilde{\bm{D}}^{-1/2}$ [47], where $\widetilde{\bm{A}}=\bm{A}+\bm{I}_{N}$ and $\widetilde{\bm{D}}=\text{diag}(\widetilde{\bm{A}}\mathbf{1}_{N})$.
+However, several influential works have also used the asymmetrically normalized adjacency, $\hat{\bm{A}}^{\text{asym}}\coloneqq\widetilde{\bm{D}}^{-1}\widetilde{\bm{A}}$ [36, 72, 50].
+2.2 DropEdge-variants
+DropEdge [68] is a random data augmentation technique that works by sampling a subgraph of the input graph in each layer, followed by the addition of self-loops, and uses that for message passing.
 Several variants of DropEdge have also been proposed, forming a family of random edge-dropping algorithms for tackling the over-smoothing problem.
-For example, DropNode (Feng et al., 2020) independently samples nodes and sets their representations to 00, followed by rescaling to make the feature matrix unbiased.
-This is equivalent to setting the corresponding columns of the propagation matrix to 00.
-In a similar vein, DropAgg (Jiang et al., 2023) samples nodes that don’t aggregate messages from their neighbors.
+For example, DropNode [23] independently samples nodes and sets their features to $\mathbf{0}$, followed by rescaling to make the feature matrix unbiased.
+This is equivalent to setting the corresponding columns of the propagation matrix to $\mathbf{0}$.
+In a similar vein, DropAgg [43] samples nodes that don’t aggregate messages from their neighbors.
 This is equivalent to dropping the corresponding rows of the adjacency matrix.
-Combining these two approaches, DropGNN (Papp et al., 2021) samples nodes which neither propagate nor aggregate messages in a given message-passing step.
+Combining these two approaches, DropGNN [65] samples nodes which neither propagate nor aggregate messages in a given layer.
 These algorithms alleviate over-smoothing by reducing the number of messages being propagated in the graph, thereby slowing down the convergence of node representations.
+2.3 Dropout-variants
+Dropout is a stochastic regularization technique which reduces over-fitting by randomly dropping features before each layer.
+It has been successful with various architectures, like CNNs [77] and transformers [80], and has also found applications in GNN training.
+DropMessage [21] is a variant of Dropout designed specifically for message-passing schemes – it acts directly on the messages over each edge, instead of the node representations.
+This reduces the induced variance in the messages compared to Dropout, DropEdge and DropNode, while at the same time making the method more effective at alleviating over-smoothing and enabling the training of deep GNNs.
 2.4 Over-squashing
-Over-squashing refers to the problem of information from exponentially growing neighborhoods Chen et al.
-(2018a) being squashed into finite-sized node representations Alon & Yahav (2021).
-This results in a loss of information as it is propagated over long distances, disallowing MPNNs from capturing long-range interactions (LRIs) and limiting their applications to short-range tasks.
+Over-squashing refers to the problem of information from exponentially growing neighborhoods [11] being squashed into finite-sized node representations [4].
 Topping et al.
-(2022) formally characterized over-squashing in terms of the Jacobian of the node-level representations w.r.t.
+[79] formally characterized over-squashing in terms of the Jacobian of the node-level representations w.r.t.
 the input features: $\|\partial\bm{z}^{\left(L\right)}_{i}/\partial\bm{x}_{j}\|_{1}$.
@@
 small perturbations in a node’s features don’t effect other distant nodes’ representations.
-Several works have linked over-squashing in an MPNN with topological properties like Cheeger’s constant Giraldo et al.
-(2023); Karhadkar et al.
-(2023), curvature of edges Topping et al.
-(2022); Nguyen et al.
-(2023); Liu et al.
-(2023), effective resistance between nodes Black et al.
-(2023); Arnaiz-Rodríguez et al.
-(2022) and the expected commute time between them Di Giovanni et al.
-(2023); Giovanni et al.
-(2024).
-These results have inspired the design of several graph rewiring techniques that strategically add edges to improve the connectivity in the graph, thereby alleviating over-squashing.
+See Appendix A for an extensive discussion of related works addressing the problems of over-smoothing and over-squashing, and a unified treatment of the two.
 3 Sensitivity Analysis
@@
 This will allow us to predict how DropEdge-variants affect communication between nodes at various distances, which is relevant for predicting their suitability towards learning LRIs.
-3.1 Linear GCNs
-We start our analysis with linear GCNs, and treat more general MPNN architectures in the following subsection. In this model, the final node representations can be summarised as
+Here, we present our analysis for linear GCNs, and treat more general nonlinear MPNN architectures in Section C.1. In this model, the final node representations can be summarised as
 $\bm{Z}^{\left(L\right)}=\left(\prod_{\ell=1}^{L}\hat{\bm{A}}^{\left(\ell\right)}\right)\bm{X}\bm{W}\in{}^{N\times H^{\left(L\right)}}$ (3.1)
@@
 where $q\in[0,1)$ is the dropping probability.
-See Section B.1 for a proof.
-Other Variants.
-We will similarly derive the expected propagation matrix for other random edge-dropping algorithms.
-First off, DropNode (Feng et al., 2020) samples nodes and drops corresponding columns from the aggregation matrix directly, followed by rescaling of its entries:
-$\begin{split}\mathbb{E}_{\mathsf{DN}}\left[\frac{1}{1-q}\hat{\bm{A}}\right]=\frac{1}{1-q}\times\left(1-q\right)\hat{\bm{A}}=\hat{\bm{A}}\end{split}$ (3.4)
-That is, the expected propagation matrix is the same as in a NoDrop model ($q=0$).
-Nodes sampled by DropAgg Jiang et al. (2023) don’t aggregate messages. Therefore, if $\hat{\bm{A}}=\hat{\bm{A}}^{\text{asym}}$, then the expected propagation matrix is given by
-$\begin{split}\mathbb{E}_{\mathsf{DA}}\left[\hat{\bm{A}}_{ii}\right]&=q+\frac{1-q}{d_{i}+1}=\frac{1+d_{i}q}{d_{i}+1}>\mathbb{E}_{\mathsf{DE}}\left[\hat{\bm{A}}_{ii}\right]\\ \mathbb{E}_{\mathsf{DA}}\left[\hat{\bm{A}}_{ij}\right]&=\frac{1}{d_{i}}\left(1-\mathbb{E}_{\mathsf{DA}}\left[\hat{\bm{A}}_{ii}\right]\right)<\mathbb{E}_{\mathsf{DE}}\left[\hat{\bm{A}}_{ij}\right]\end{split}$ (3.5)
-Finally, DropGNN (Papp et al., 2021) samples nodes which neither propagate nor aggregate messages.
-From any node’s perspective, if it is not sampled, then its aggregation weights are computed as for DropEdge:
-$\begin{split}\mathbb{E}_{\mathsf{DG}}\left[\hat{\bm{A}}_{ii}\right]&=q+\left(1-q\right)\mathbb{E}_{\mathsf{DE}}\left[\hat{\bm{A}}_{ii}\right]=q+\frac{1-q^{d_{i}+1}}{d_{i}+1}>\mathbb{E}_{\mathsf{DA}}\left[\hat{\bm{A}}_{ii}\right]\\ \mathbb{E}_{\mathsf{DG}}\left[\hat{\bm{A}}_{ij}\right]&=\frac{1}{d_{i}}\left(1-\mathbb{E}_{\mathsf{DG}}\left[\hat{\bm{A}}_{ii}\right]\right)<\mathbb{E}_{\mathsf{DA}}\left[\hat{\bm{A}}_{ij}\right]\end{split}$ (3.6)
+See Section B.1 for a proof, and a similar treatment of DropNode, DropAgg and DropGNN.
 1-Layer Linear GCNs. $\forall q\in\left(0,1\right)$ we have
 $\begin{split}\dot{\bm{P}}_{ii}&=\frac{1}{d_{i}+1}\sum_{k=0}^{d_{i}}q^{k}>\frac{1}{d_{i}+1}\\ \dot{\bm{P}}_{ij}&=\frac{1}{d_{i}}\left(1-\dot{\bm{P}}_{ii}\right)<\frac{1}{d_{i}+1}\end{split}$ (3.4)
-where the right-hand sides of the two inequalities are the corresponding entries in the propagation matrix of a NoDrop model. Equation 3.3 to Equation 3.7 together imply the following result:
+where the right-hand sides of the two inequalities are the corresponding entries in the propagation matrix of a NoDrop model. s 3.3, 3.4, B.16 and B.17 together imply the following result:
 Lemma 3.2.
@@
     decreases the sensitivity to its neighbors’ features.
-In other words, DropEdge-variants prevent a 1-layer GCN from fully utilizing neighborhood information when learning node representations.
-Ignoring the graph topology this way makes the model resemble an MLP, limiting its expressiveness and hindering its ability to model graph-data.
-Figure 1: Entries of $\dot{\bm{P}}^{6}$, averaged over molecular graphs sampled from the Proteins dataset.
-Left: Sensitivity between nodes decays at a polynomial rate w.r.t.
-their distance.
-Middle: Similarly, it decays at a polynomial rate w.r.t.
-the DropEdge probability.
-Right: It decays more quickly with DropAgg than with DropEdge, and even more rapidly with DropGNN.
+(a) Entries of $\dot{\bm{P}}^{6}$ decay at exponential rate w.r.t. distance between nodes, and polynomial rate w.r.t. to the DropEdge probability.
+(b) MC-approximation of influence distribution in ReLU-GCNs.
+Figure 1: Empirical sensitivity analysis using the Cora dataset.
 L-layer Linear GCNs.
@@
 For nodes L-hops away, however, we can show that DropEdge always decreases the corresponding entry in $\dot{\bm{P}}^{L}$, reducing the effective reachability of GCNs.
-Using Equation 3.5 and Equation 3.6, we can show the same for DropAgg and DropGNN, respectively.
+Using s B.16 and B.17, we can show the same for DropAgg and DropGNN, respectively.
 Theorem 3.1.
 In an L-layer linear GCN with $\hat{\bm{A}}=\hat{\bm{A}}^{\text{asym}}$, using DropEdge, DropAgg or DropGNN decreases the sensitivity of a node $i\in\mathcal{V}$ to another node $j\in\mathbb{S}^{\left(L\right)}\left(i\right)$, thereby reducing its effective receptive field.
-Moreover, the sensitivity monotonically decreases as the dropping probability is increased.
+Moreover, the sensitivity decreases with increasing dropping probability.
 See Section B.2 for a precise quantitative statement and the proof.
@@
 Although no general statement could be made about the change in sensitivity between nodes up to $L-1$ hops away, we can analyze such pairs empirically.
-We sampled 100 molecular graphs from the Proteins dataset (Dobson & Doig, 2003), binned the node-pairs in each graph by the shortest distance between them, and then plotted the average of the corresponding entries in $\dot{\bm{P}}^{L}$, $L=6$.
-The results are shown in Figure 1.
-In the left subfigure, we observe that the sensitivity between two nodes decays at a polynomial rate with increasing distance between them.
-Moreover, DropEdge increases the expected sensitivity between nodes close to each other (0-hop and 1-hop neighbors) in the original topology, but reduces it between nodes farther off.
-In the middle subfigure, we show how the average sensitivity at different distances changes with the DropEdge probability.
-Specifically, we observe that the decay in sensitivity to nodes at large distances is polynomial in the DropEdge probability, which suggests that the algorithm would not be suitable for capturing LRIs.
+We compute the L-hop transition matrix $\dot{\bm{P}}^{L}$ – proportional to expected sensitivity in linear GCNs under DropEdge – for the Cora dataset, and average the entries after binning node pairs by the shortest distance between them.
+The results are shown in 1(a).
+In the left subfigure, we observe that the expected sensitivity decays at an exponential rate with increasing distance between the corresponding nodes.
+In the middle subfigure, we observe that DropEdge increases the expected sensitivity between nodes close to each other (0-hop and 1-hop neighbors) in the original topology, but reduces it between nodes farther off.
 Similar conclusions can be made with the symmetrically normalized propagation matrix (see Section D.1).
-Finally, in the right subfigure, we compare the DropEdge-variants by plotting the sensitivity at $d_{\mathcal{G}}\left(i,j\right)=L=6$.
-Although the analytical form of all the expected propagation matrices are available to us, we approximate them using Monte-Carlo sampling.
-As expected from Equation 3.5, we observe that DropAgg not only decreases the sensitivity of node representations to distant nodes, but does it to a greater extent than DropEdge.
-Similarly, Equation 3.6 suggested that DropGNN could be even more harmful than DropAgg, and this is validated by the empirical results.
-In fact, for $q>0.7$, nodes 6-hops away are mostly insensitive to each other.
-3.2 Nonlinear MPNNs
+Note that the over-squashing effects of DropAgg and DropGNN would, in theory, be even more severe, as suggested by s B.16 and B.17.
+Nonlinear MPNNs.
 While linear networks are useful in simplifying the theoretical analysis, they are often not practical.
-In this subsection, we will consider the upper bounds on sensitivity established in previous works, and extend them to the DropEdge setting.
-ReLU GCNs.
-Xu et al.
-(2018) considered the case of ReLU nonlinearity, so that the update rule is $\bm{Z}^{\left(\ell\right)}=\mathsf{ReLU}(\hat{\bm{A}}\bm{Z}^{\left(\ell-1\right)}\bm{W}^{\left(\ell\right)})$.
-Additionally, it makes the simplifying assumption that each path in the computational graph is active with a fixed probability, $\rho$ (Kawaguchi, 2016, Assumption A1p-m).
-Accordingly, the sensitivity (in expectation) between any two nodes is given as
-$\left\|\mathbb{E}_{\mathsf{ReLU}}\left[\frac{\partial\bm{z}^{\left(L\right)}_{i}}{\partial\bm{x}_{j}}\right]\right\|_{1}=\left[\rho\left\|\prod_{\ell=1}^{L}\bm{W}^{\left(\ell\right)}\right\|_{1}\right]\left(\hat{\bm{A}}^{L}\right)_{ij}=\zeta_{1}^{\left(L\right)}\left(\hat{\bm{A}}^{L}\right)_{ij}$ (3.8)
-where $\zeta_{1}^{\left(L\right)}$ is independent of the choice of nodes $i,j\in\mathcal{V}$. Taking an expectation w.r.t. the random edge masks, we get
-$\mathbb{E}_{\mathbf{M}^{\left(1\right)},\ldots,\mathbf{M}^{\left(L\right)}}\left[\left\|\mathbb{E}_{\mathsf{ReLU}}\left[\frac{\partial\bm{z}^{\left(L\right)}_{i}}{\partial\bm{x}_{j}}\right]\right\|_{1}\right]$ $=\zeta_{1}^{\left(L\right)}\mathbb{E}_{\mathbf{M}^{\left(1\right)},\ldots,\mathbf{M}^{\left(L\right)}}\left[\left(\prod_{\ell=1}^{L}\hat{\bm{A}}^{\left(\ell\right)}\right)_{ij}\right]$ (3.9)
-$=\zeta_{1}^{\left(L\right)}\left(\mathbb{E}\left[\hat{\bm{A}}\right]^{L}\right)_{ij}$ (3.10)
-Using Theorem 3.1, we conclude that in a ReLU-GCN, DropEdge, DropAgg and DropGNN will reduce the expected sensitivity between nodes L-hops away.
-Empirical observations in Figure 1 and Figure 4 suggest that we may expect an increase in sensitivity to neighboring nodes, but a significant decrease in sensitivity to those farther away.
-Source-only Message Functions. Black et al. (2023, Lemma 3.2) considers MPNNs with aggregation functions of the form
-$\mathsf{Agg}^{\left(\ell\right)}\left(\bm{z}_{i}^{\left(\ell-1\right)},\left\{\bm{z}_{j}^{\left(\ell-1\right)}:j\in\mathbb{S}^{\left(1\right)}\left(i\right)\right\}\right)=\sum_{j\in\mathbb{B}^{\left(1\right)}\left(i\right)}\hat{\bm{A}}_{ij}\mathsf{Msg}^{\left(\ell\right)}\left(\bm{z}_{j}^{\left(\ell-1\right)}\right)$ (3.11)
-and $\mathsf{Upd}$ and $\mathsf{Msg}$ functions with bounded gradients. In this case, the sensitivity between two nodes $i,j\in\mathcal{V}$ can be bounded as
-$\left\|\frac{\partial\bm{z}^{\left(L\right)}_{i}}{\partial\bm{x}_{j}}\right\|_{1}\leq\zeta_{2}^{\left(L\right)}\left(\sum_{\ell=0}^{L}\hat{\bm{A}}^{\ell}\right)_{ij}$ (3.12)
-As before, we can use the independence of edge masks to get an upper bound on the expected sensitivity:
-$\mathbb{E}_{\mathbf{M}^{\left(1\right)},\ldots,\mathbf{M}^{\left(L\right)}}\left[\left\|\frac{\partial\bm{z}^{\left(L\right)}_{i}}{\partial\bm{x}_{j}}\right\|_{1}\right]$ $\leq\zeta_{2}^{\left(L\right)}\left(\mathbb{E}_{\mathbf{M}^{\left(1\right)},\ldots,\mathbf{M}^{\left(L\right)}}\left[I_{N}+\sum_{\ell=1}^{L}\prod_{k=1}^{\ell}\hat{\bm{A}}^{\left(k\right)}\right]\right)_{ij}$ (3.13)
-$=\zeta_{2}^{\left(L\right)}\left(\sum_{\ell=0}^{L}\mathbb{E}\left[\hat{\bm{A}}\right]^{\ell}\right)_{ij}$ (3.14)
-Figure 5 shows the plot of the entries of $\sum_{\ell=0}^{6}\dot{\bm{P}}^{\ell}$ (i.e.
-for DropEdge), as in the upper bound above, with $\hat{\bm{A}}=\hat{\bm{A}}^{\text{asym}}$.
-We observe that the sensitivity between nearby nodes marginally increases, while that between distant nodes notably decreases (similar to Figure 1), suggesting significant over-squashing.
-Similar observations can be made with $\hat{\bm{A}}=\hat{\bm{A}}^{\text{sym}}$, and for other DropEdge-variants.
-Source-and-Target Message Functions. Topping et al. (2022, Lemma 1) showed that if the aggregation function is instead given by
-$\mathsf{Agg}^{\left(\ell\right)}\left(\bm{z}_{i}^{\left(\ell-1\right)},\left\{\bm{z}_{j}^{\left(\ell-1\right)}:j\in\mathbb{S}^{\left(1\right)}\left(i\right)\right\}\right)=\sum_{j\in\mathbb{B}^{\left(1\right)}\left(i\right)}\hat{\bm{A}}_{ij}\mathsf{Msg}^{\left(\ell\right)}\left(\bm{z}_{i}^{\left(\ell-1\right)},\bm{z}_{j}^{\left(\ell-1\right)}\right)$ (3.15)
-then the sensitivity between nodes $i\in\mathcal{V}$ and $j\in\mathbb{S}^{\left(L\right)}\left(i\right)$ can be bounded as
-$\left\|\frac{\partial\bm{z}^{\left(L\right)}_{i}}{\partial\bm{x}_{j}}\right\|_{1}\leq\zeta_{3}^{\left(L\right)}\left(\hat{\bm{A}}^{L}\right)_{ij}$ (3.16)
-With random edge-dropping, this bound can be adapted as follows:
-$\mathbb{E}_{\mathbf{M}^{\left(1\right)},\ldots,\mathbf{M}^{\left(L\right)}}\left[\left\|\frac{\partial\bm{z}^{\left(L\right)}_{i}}{\partial\bm{x}_{j}}\right\|_{1}\right]\leq\zeta_{3}^{\left(L\right)}\left(\mathbb{E}\left[\hat{\bm{A}}\right]^{L}\right)_{ij}$ (3.17)
-which is similar to Equation 3.10, only with a different proportionality constant, that is anyway independent of the choice of nodes.
-Here, again, we invoke Theorem 3.1 to conclude that $(\mathbb{E}[\hat{\bm{A}}]^{L})_{ij}$ decreases monotonically with increasing DropEdge probability $q$.
-This implies that, in a non-linear MPNN with $\hat{\bm{A}}=\hat{\bm{A}}^{\text{asym}}$, DropEdge lowers the sensitivity bound given above.
-Empirical results in Figure 4 support the same conclusion for $\hat{\bm{A}}=\hat{\bm{A}}^{\text{sym}}$.
-Message of the Section: Studying the expected propagation matrix allows us to predict the effect of random edge-dropping methods on information propagation in MPNNs.
-Specifically, DropEdge, DropAgg and DropGNN increase the sensitivity of nodes to their neighbors, but decrease it to nodes farther off.
-This suggests that these methods would be unsuitable for long-range tasks, where it is imperative to facilitate communication between distant nodes.
+In Section C.1, we treat the upper bounds on sensitivity established in previous works, and extend them to the DropEdge-variants.
+Even still, although theoretical bounds offer valuable guarantees, they can be arbitrarily loose in the absence of error quantification, making their practical relevance unclear.
+To reliably conclude the empirical behaviour of DropEdge- and Dropout-variants, we turn to Monte Carlo simulations with ReLU-GCNs; see Section D.3 for a description of the experiment setup.
+1(b) compares the influence of the source nodes [84] at different distances using a dropout probability of $0.5$.
+We observe that while the effect of DropNode on the sensitivity profile – as compared to the baseline NoDrop – is relatively insignificant, models using DropEdge, DropAgg and DropGNN have remarkably lower sensitivity to distant nodes, as predicted by our theory.
+4 Sensitivity-Aware DropEdge
+Lemma 3.1 tells us that DropEdge decreases the weight of cross-edges, $\left(j\to i\right)$, in the expected propagation matrix, i.e.
+the strength of message passing over these edges decreases.
+The fraction of information preserved over a cross-edge is dependent only on the dropping probability and the target node’s in-degree, $d_{i}$.
+We can directly control this quantity using a per-edge dropping probability, $q_{i}$, dependent only on the receiving node’s in-degree:
+$\begin{split}c&=\frac{\mathbb{E}_{\mathsf{DE}}[\hat{\bm{A}}_{ij}]}{\mathbb{E}_{\mathsf{ND}}[\hat{\bm{A}}_{ij}]}=\frac{d_{i}+1}{d_{i}}\left(1-\frac{1-q_{i}^{d_{i}+1}}{\left(1-q_{i}\right)\left(d_{i}+1\right)}\right)\Longrightarrow 1-c=\frac{q_{i}-q_{i}^{d_{i}+1}}{d_{i}\left(1-q_{i}\right)}\end{split}$ (4.1)
+where $c$ is the fraction of information preserved, e.g.
+95%.
+We can solve for $q_{i}$ and mask the incoming edges to node $i$ accordingly; we name this algorithm DropSens.
+In Section E.3, we present a Python implementation of the algorithm, as well as a computationally efficient approximation to Equation 4.1.
+In 1(b), we observe that DropSens improves sensitivity between distant nodes, compared to DropEdge.
 5 Experiments
-Our theoretical analysis indicates that DropEdge may degrade the performance of GNNs on tasks that depend on capturing LRIs.
-In this section, we test this hypothesis by evaluating DropEdge models on both short-range and long-range tasks.
-4.1 Setup
-Datasets.
-Although identifying whether a task requires modeling LRIs can be challenging, understanding the structure of the datasets can provide some insight.
+Our theoretical analysis indicates that random dropping may degrade the performance of GNNs in tasks that depend on capturing LRIs.
+In this section, we test this hypothesis by evaluating DropEdge- and Dropout-variants on both synthetic and real-world datasets.
+A complete description of the datasets is provided in Section E.1, and the experimental details are in Section E.2.
+5.1 Synthetic Datasets
+Figure 2: Train and test MAE of 11-layer GCNs on the SyntheticZINC dataset, averaged over 10 initializations.
+SyntheticZINC [33] is a synthetic variant of the ZINC dataset [42], designed to study the effect of information mixing in graph learning.
+Node features are sparsely assigned, and the target requires non-linear mixing of two selected nodes’ features, chosen based on their commute time [10].
+We vary the mixing level and evaluate an 11-layer GCN, ensuring sufficient message passing.
+For better readability, we only test DropEdge, Dropout and DropMessage – the three more popular methods used for training deep GNNs.
+The results are presented in Figure 2, where we can observe that the mean absolute error (MAE) increases with the commute time percentile used to select the node pairs, as was hypothesized and evidenced in Giovanni et al.
+[33].
+Additionally, we observe that both train and test performance decline when using dropout with a probability as low as 0.2, and even more so with a higher probability of 0.5.
+These results provide strong evidence for the detrimental effects of dropout methods in modelling long-range interactions, supporting our theoretical analysis.
+5.2 Real-world Datasets
+Table 1: Difference in mean test accuracy (%) between the best performing configuration of each dropout method and the baseline NoDrop model.
+Cell colors represent p-values from a t-test evaluating whether dropout improves performance: green indicates significance at 90% confidence, while red denotes insignificant results.
+(a) Node classification tasks.
+GNN Dropout Homophilic Networks Heterophilic Networks
+Cora CiteSeer PubMed Chameleon Squirrel TwitchDE
+GCN DropEdge $+0.419$ $+0.686$ $+0.385$ $-0.634$ $+0.009$ $-0.093$
+DropNode $+0.373$ $+0.197$ $+0.841$ $-0.674$ $-0.656$ $-0.113$
+DropAgg $+0.224$ $+0.486$ $-0.245$ $-10.640$ $-13.970$ $-6.674$
+DropGNN $+0.396$ $+0.820$ $+0.506$ $-1.774$ $-0.291$ $-0.395$
+Dropout $+0.628$ $+0.128$ $+1.218$ $-1.395$ $-0.110$ $-0.260$
+DropMessage $+0.030$ $-0.388$ $+1.217$ $+1.322$ $+0.320$ $+0.160$
+GIN DropEdge $-0.276$ $-0.529$ $-0.034$ $-0.726$ $+0.289$ $-0.124$
+DropNode $+0.379$ $+0.926$ $+0.296$ $-1.832$ $-0.265$ $-0.171$
+DropAgg $+0.095$ $-0.254$ $-0.487$ $-1.153$ $+0.275$ $+0.205$
+DropGNN $-0.478$ $-2.284$ $-1.366$ $-1.642$ $-0.066$ $-0.008$
+Dropout $+1.347$ $+0.011$ $+0.368$ $-2.702$ $-0.152$ $-0.573$
+DropMessage $+2.602$ $+0.219$ $+1.030$ $+0.943$ $+0.108$ $-0.025$
+(b) Graph classification tasks.
+GNN Dropout Molecular Networks Social Networks
+Mutag Proteins Enzymes Reddit IMDb Collab
+GCN DropEdge $-1.100$ $+1.750$ $-0.589$ $-8.380$ $+1.300$ $-1.145$
+DropNode $-5.100$ $+2.339$ $-2.292$ $-7.440$ $+2.720$ $-4.054$
+DropAgg $-0.900$ $+1.214$ $-0.553$ $-16.460$ $+3.580$ $-30.386$
+DropGNN $-0.200$ $+1.589$ $-2.918$ $-11.860$ $+1.540$ $+0.705$
+Dropout $-0.300$ $+2.018$ $-6.208$ $-6.050$ $+1.240$ $-1.729$
+DropMessage $+2.000$ $+2.143$ $-4.906$ $-7.360$ $+1.300$ $-0.330$
+GIN DropEdge $-1.100$ $-1.804$ $-4.716$ $-2.700$ $-1.820$ $-0.642$
+DropNode $-3.800$ $-2.911$ $-0.493$ $+0.550$ $-0.120$ $+1.206$
+DropAgg $-3.800$ $-1.982$ $-0.640$ $+0.930$ $-1.640$ $-5.943$
+DropGNN $-7.000$ $-2.750$ $-3.694$ $+1.210$ $-5.200$ $-6.593$
+Dropout $-2.400$ $-3.446$ $-2.047$ $+2.590$ $-1.180$ $-0.180$
+DropMessage $-3.600$ $-1.125$ $-0.302$ $+0.850$ $-0.460$ $+1.126$
+To test the dropping methods on real-world datasets, we use the GCN, GIN [85] and GAT [81] architectures – GCN and GIN satisfy the model assumptions made in all the theoretical results presented in Section 3, while GAT does not satisfy any of them, since the attention scores are computed as a function of all the node representations.
+Therefore, GCN, GIN and GAT together provide a broad representation of different MPNN architectures.
+We present the results for GCN and GIN in the main text, since these models were used as baselines in a majority of works on alleviating over-squashing [8, 44, 79, 35, 4, 5, 67]; the results for GAT are reported in Table 6.
+For each dataset--model--dropout combination, we perform 20 independent runs to find the best performing dropout configuration; results are reported in Table 9.
+We then perform a t-test to assess whether dropout improves performance, using 50 samples from the NoDrop model ($q=0$) and 50 samples from the best performing dropout configuration.
+[^3] The t-test assumes that both samples are drawn from normal distributions – all Shapiro-Wilk tests for non-normality of samples [74] failed at 90% confidence.
+In this section, we report the p-values of the tests, and in Table 7, we report the effect sizes as Hedges’ $g$ statistic [39].
+Node-classification.
+Although determining whether a task requires modelling LRIs can be challenging, understanding the structure of the datasets can provide important insight.
 For example, homophilic datasets have local consistency in node labels, i.e.
@@
 On the other hand, in heterophilic datasets, nearby nodes often have dissimilar labels.
-Since DropEdge-variants increase the sensitivity of a node’s representations to its immediate neighbors, and reduce its sensitivity to distant nodes, we expect it to improve performance on homophilic datasets but harm performance on heterophilic ones (see Topping et al.
-(2022)).
-In this work, we use Cora (McCallum et al., 2000) and CiteSeer (Giles et al., 1998) as representatives of homophilic datasets (Zhu et al., 2020), and Squirrel, Chameleon and TwitchDE (Rozemberczki et al., 2021) to represent heterophilic datasets (Lim et al., 2021).
-The networks’ statistics are presented in Table 1, where we can note the significantly lower homophily measures of heterophilic datasets.
-Dataset Nodes Edges Features Classes Homophily
-Cora 2,708 10,556 1,433 7 0.766
-CiteSeer 3,327 9,104 3,703 6 0.627
-Chameleon 2,277 36,051 2,325 5 0.062
-Squirrel 5,201 216,933 2,089 5 0.025
-TwitchDE 9,498 306,276 128 2 0.142
-Table 1: Dataset statistics. Number of edges excludes self-loops. Homophily measures from Lim et al. (2021).
-Models.
-We use two MPNN architectures, GCN and GAT, to demonstrate the effect of DropEdge.
-GCN satisfies the model assumptions made in all the theoretical results presented in Section 3, while GAT does not fit into any of them.
-Therefore, GCN and GAT together provide a broad representation of different MPNN architectures.
-For GAT, we use 2 attentions heads in order to keep the computational load manageable, while at the same time harnessing the expressiveness of the multi-headed self-attention mechanism.
-As for the model depth, we vary it as $L=2,4,6,8$.
-Dropout Methods.
-In addition to the DropEdge and its variants – DropNode, DropAgg and DropGNN – discussed in Section 3, we evaluate related stochastic regularization methods like regular dropout (Srivastava et al., 2014) (referred as Dropout), DropMessage (Fang et al., 2023) and SkipNode Lu et al.
-(2024).
-This allows us to study the pervasiveness of the problem of over-squashing over a large set of methods designed for alleviating over-smoothing.
-Dropping Probability.
-For all the methods, the dropping probabilities are varied as $q=0.0,0.1,\ldots,0.9$, so as to reliably capture the trends in model accuracy.
-We adopt the common practice of turning the methods off at test-time ($q=0$), isolating their effect on optimization and generalization, which our theory does not address.
-Other experimental details are reported in Appendix E.
-We conduct 5 independent runs for each dataset--model--dropout--probability configuration and report the average test accuracy (with early-stopping using the validation set).
-4.2 Results
-Dropout GNN Homophilic Heterophilic
-Cora CiteSeer Chameleon Squirrel TwitchDE
-DropEdge GCN $+0.396$ $+0.448$ $-0.723$ $-0.614$ $-0.509$
-GAT $+0.547$ $+0.548$ $-0.236$ $-0.641$ $-0.390$
-DropNode GCN $-0.320$ $-0.662$ $-0.863$ $-0.790$ $-0.465$
-GAT $-0.706$ $-0.735$ $-0.855$ $-0.286$ $-0.512$
-DropAgg GCN $-0.015$ $+0.294$ $-0.334$ $-0.317$ $-0.566$
-GAT $-0.019$ $+0.392$ $-0.353$ $-0.398$ $-0.022$
-DropGNN GCN $+0.360$ $+0.468$ $-0.746$ $-0.600$ $-0.671$
-GAT $+0.236$ $+0.525$ $-0.507$ $-0.304$ $-0.139$
-Dropout GCN $+0.034$ $-0.309$ $-0.610$ $-0.465$ $-0.300$
-GAT $+0.120$ $-0.110$ $-0.836$ $-0.545$ $-0.397$
-DropMessage GCN $-0.220$ $-0.607$ $-0.404$ $-0.364$ $-0.199$
-GAT $+0.028$ $-0.368$ $-0.530$ $-0.697$ $-0.148$
-SkipNode GCN $+0.161$ $+0.017$ $-0.528$ $-0.285$ $-0.244$
-GAT $+0.403$ $+0.466$ $-0.436$ $+0.248$ $+0.278$
-Table 2: Spearman correlation between dropping probability and test accuracy.
-Note the positive correlations for homophilic datasets and negative correlations for heterophilic datasets.
-Results suggest that random edge-dropping, although effective at improving generalization in short-range tasks, is unsuitable for long-range ones.
-In Table 2, we report the rank correlation between the dropping probability and test accuracy of different dataset--model--dropout combinations.
-The statistics are computed over $5010\ (q=0.0,0.1,\ldots,0.9)\times 5\ (\text{samples})=$ 50 runs for each $L=2,4,6,8$, and then averaged.
-It is clear to see that in most combinations with the homophilic datasets Cora and CiteSeer, the correlation is positive, indicating that DropEdge and its variants improve test-time performance at short-range tasks.
-On the other hand, with the heterophilic datasets Chameleon, Squirrel and TwitchDE, the correlation values are negative.
-This suggests that the edge-dropping methods harm generalization in long-range tasks by forcing the model to overfit to short-range signals.
-(a) Homophilic datasets
-(b) Heterophilic datasets
-Figure 2: Dropping probability versus test accuracy of DropEdge-GCN.
-The theory the explains the contrasting trends as follows: random edge-dropping pushes models to fit to local information during training, which is suitable for short-range tasks, but harms test-time performance in long-range ones.
-Figure 2 allows us to visualize these trends for the homophilic and heterophilic datasets.
-For conciseness, we only display the results for DropEdge-GCNs.
-Clearly, on Cora and CiteSeer, the model’s test-time performance improves with increasing dropping probability, as has been reported earlier.
-However, on Chameleon, Squirrel and TwitchDE, the performance degrades with increasing dropping probability, as was suggested by our theoretical results.
-This highlights an important gap in our understanding of dropout methods – while their positive effects on model performance have been well-studied, making them popular choices for training deep GNNs, their evaluation has been limited to short-range tasks, leaving their negative impact on capturing LRIs overlooked.
-Remark on DropNode.
-In Equation 3.4, we noted that DropNode does not suffer from loss in sensitivity.
-However, note that those results were in expectation.
-Moreover, our analysis did not account for the effects on the learning trajectory.
-In practice, a high DropNode probability would make it hard for information in the node features to reach distant nodes.
-This would prevent the model from learning to effectively combine information from large neighborhoods, harming generalization.
-In fact, in Table 2, we can see that it is the only dropping method with a negative correlation between dropping probability and test accuracy for each dataset--model combination, including homophilic ones.
-See Figure 6 for a visualization of its negative effects on test accuracy.
-4.3 Over-squashing or Under-fitting?
-Figure 3: DropEdge probability versus training accuracy of GCNs.
-The training performance improves with $q$, suggesting that the models are not underfitting.
-Instead, the reason for poor test-time performance (2(a)) is that models are over-fitting to short-range signals during training, resulting in poor generalization.
-The results in the previous subsection suggest that using random edge-dropping to regularize model training leads to poor test-time performance.
-We hypothesize that this occurs because the models struggle to propagate information over long distances, causing overfitting of node representations to local neighborhoods.
-However, a confounding factor is at play: DropEdge variants reduce the generalization gap by preventing overfitting to the training set, which manifests as reduced training performance.
-If this regularization is too strong, it could lead to underfitting, which could also explain the poor test-time performance on heterophilic datasets.
-This concern is particularly relevant because the heterophilic networks are much larger than homophilic ones (Table 1), making them more prone to underfitting.
-To investigate this, we plot the training accuracies of deep DropEdge-GCNs on the heterophilic datasets; Figure 3 shows the results.
-It is clear that the models do not underfit as the dropping probability increases.
-In fact, somewhat unexpectedly, the training metrics improve.
-Together with the results in Figure 2, we conclude that DropEdge-like methods are detrimental in long-range tasks, as they cause overfitting to short-range artifacts in the training data, resulting in poor generalization at test-time.
-Message of the Section: Random edge-dropping algorithms reduce the receptive field of MPNNs, forcing them to fit to short-range signals in the training set.
-While this may make deep GNNs suitable for homophilic datasets, it results in overfitting on heterophilic datasets, which leads to poor test-time performance.
-Therefore, these methods should be used only after carefully understanding the task at hand.
+Since DropEdge-variants increase the sensitivity of a node’s representations to its immediate neighbors, and reduce its sensitivity to distant nodes, we expect it to improve performance on homophilic datasets but harm performance on heterophilic ones; such a setup was also used in Topping et al.
+[79].
+In this work, we use Cora [57], CiteSeer [31] and PubMed [62] as representatives of homophilic datasets [98, 52], and Squirrel, Chameleon and TwitchDE [69] to represent heterophilic datasets [52].
+The networks’ statistics are presented in Table 4, where we can note the remarkably lower homophily measures of heterophilic datasets.
+The results are presented in 1(a).
+It is clear to see that dropout significantly improves test performance on homophilic datasets – with $40/54\approx 74\%$ cases performing better than the corresponding NoDrop baseline – indicating that these methods are indeed beneficial in tackling short-range tasks.
+On the other hand, with the heterophilic datasets, the improvement is insignificant.
+Rather, in most ($45/54\approx 83\%$) cases, the best dropout configuration performs worse than the NoDrop baseline.
+This suggests that the dropping methods harm generalization in long-range tasks by forcing models to overfit to short-range signals (see Section F.3 for supporting evidence).
+Graph-classification.
+Several graph classification datasets have also been identified as long-range tasks, like the molecular networks datasets Mutag [17], Proteins [19] and Enzymes [9], and the social networks datasets Reddit, IMDb and Collab [87].
+These datasets have also been used for evaluation in previous works on over-squashing, including Karhadkar et al.
+[44], Black et al.
+[8].
+The results are shown in 1(b), where we observe that dropout methods generally have insignificant effects on model performance, and often even a non-positive effect ($67/108\approx 62\%$ cases).
+Notably, the p-values are lower as compared to those recorded for heterophilic datasets in 1(a), i.e.
+higher evidence for efficacy of dropping methods.
+We conjecture that over-squashing may have limited impact on model performance in graph-level tasks since the aggregation module eventually mixes information from distant nodes for computing graph-level representations.
+5.3 Evaluating DropSens
+Figure 3: Relative change in test-time performance of a GCN using DropSens, compared to the baseline DropEdge, on real-world datasets from Section 5.2.
+We start by comparing DropSens with DropEdge on real-world datasets from Section 5.2, illustrating how algorithms can be readily adapted for better suitability at modelling LRIs.
+In Figure 3, we present the relative change in error rate ($1-\text{Acc}$) of DropSens w.r.t.
+DropEdge, with GCN as the base model.
+It is clear to observe a uniform improvement in the performance on long-range tasks, suggesting that addressing over-squashing using DropSens can enhance the effectiveness of GCNs.
+Table 2: Performance of GCN with graph rewiring methods. First, second, and third best results are coloured.
+(a) Node-classification tasks – results for other methods taken from [79, Table 2].
+Rewiring Cora CiteSeer PubMed Chameleon Squirrel Actor
+None $81.89$ $72.31$ $78.16$ $41.33$ $30.32$ $23.84$
+Undirected - - - $42.02$ $35.53$ $21.45$
++FA $81.65$ $70.47$ 79.48 $42.67$ $36.86$ $24.14$
+DIGL (PPR) 83.21 73.29 $78.84$ $42.02$ $33.22$ $24.77$
+DIGL + Undirected - - - $42.68$ $32.48$ 25.45
+SDRF 82.76 72.58 79.10 42.73 37.05 28.42
+SDRF + Undirected - - - 44.46 37.67 28.35
+DropSens 84.98 73.35 84.30 53.01 41.32 $22.38$
+(b) Graph-classification tasks – results for other methods from [8, Table 1].
+Rewiring Mutag Proteins Enzymes Reddit IMDb Collab
+None $72.15$ $70.98$ $27.67$ $68.26$ 49.77 $33.78$
+Last FA $70.05$ 71.02 $26.47$ $68.49$ $48.98$ $33.32$
+Every FA $70.45$ $60.04$ $18.33$ $48.49$ $48.17$ 51.80
+DIGL 79.70 $70.76$ 35.72 76.04 64.39 54.50
+SDRF $71.05$ $70.92$ 28.37 $68.62$ $49.40$ $33.45$
+FoSR 80.00 73.42 $25.07$ 70.33 $49.66$ $33.84$
+GTR 79.10 72.59 $27.52$ $68.99$ 49.92 $33.05$
+DropSens $70.20$ $70.61$ 36.01 77.44 $49.32$ 61.68
+We now benchmark DropSens against state-of-the-art graph-rewiring techniques designed specifically to tackle over-squashing (see Section A.3 for their descriptions).
+We train a GCN on node classification tasks, following the setup in [79], and both GCN and GIN on graph classification tasks, following Karhadkar et al.
+[44], Black et al.
+[8].
+The results with GCN are reported in 2(a) and 2(b), where we find that DropSens outperforms other methods in node classification tasks, and performs competitively in graph classification tasks.
+In addition to superior performance, another advantage of DropSens over the other methods is that it significantly reduces the number of messages being propagated, thereby tackling the problem of over-smoothing and increasing training speed.
+The results with GIN are presented in Section G.3, where we observe that DropSens does not perform competitively – unsurprising, since DropSens was specifically designed to work with GCN’s message-passing scheme.
 6 Conclusion
-Our analysis points out a key assumption in algorithms designed for training deep GNNs: the idea that if a deep GNN is trainable, it must be able to model LRIs.
-Our results suggest that this, in fact, need not be true – we theoretically and empirically show that DropEdge-like algorithms exacerbate the over-squashing problem in deep GNNs, and degrade their performance on long-range tasks.
-Our results highlight a need for a thorough evaluation of methods employed when training deep GNNs, with regards to their capacity to capture LRIs.
-This will allow us to reliably deploy them in real-life, since we can be assured that the models did not simply overfit to short-range signals.
+There exists an important gap in our understanding of several algorithms designed for training deep GNNs – while their positive effects on model performance have been well-studied, making them popular choices for training deep GNNs, their evaluation has been limited to short-range tasks.
+This is rooted in a key assumption: that if a deep GNN is trainable, it must also be capable of modelling LRIs.
+As a result, potential adverse effects of these algorithms on capturing LRIs have been overlooked.
+Our results challenge this assumption – we theoretically and empirically show that DropEdge- and Dropout-variants exacerbate the over-squashing problem in deep GNNs, and degrade performance on long-range tasks.
+This highlights the need for a more comprehensive evaluation of common training practices for deep GNNs, with special emphasis on their capacity to capture LRIs.
+This is crucial for building confidence in their use beyond controlled benchmarks.
 Limitations.
-While our theoretical analysis successfully predicts how DropEdge-variants affect test performance on short-range and long-range tasks, it is based on several simplifying assumptions.
-These assumptions, although standard in the literature, limit the generalizability of our conclusions to other architectures.
-Specifically, our analysis focuses on certain classes of MPNNs, excluding several GNN architectures specifically designed to enhance long-distance information propagation (see Singh (2024) for a review).
-Practical Considerations.
-Previous studies have shown that random edge-dropping algorithms can effectively enhance generalization performance in short-range tasks, and our findings support this conclusion.
-However, we have also demonstrated that such algorithms can negatively impact over-squashing in MPNNs and harm test-time performance in long-range tasks.
-Therefore, we recommend exercising caution when using such methods, as careless application can result in models that generalize poorly, which can be detrimental in critical applications.
+While our theoretical analysis successfully predicts how DropEdge-variants affect test performance on short-range and long-range tasks, it is based on several simplifying assumptions on the message-passing scheme.
+These assumptions, although standard in the literature, limit the generalizability of our conclusions to other architectures, including ResGCNs [51], GATs [81], and Graph Transformers [83].
+Additionally, an important limitation of DropSens is that it requires an architecture-specific alteration to the edge-dropping strategy, which is not practical in general.
+As also mentioned in Section 4, we did not intend to introduce DropSens as a benchmark, but rather to demonstrate how methods designed for alleviating over-smoothing can be readily adapted to simultaneously control over-squashing.
 Future Directions.
-This work focuses on node-classification tasks, but it is also important to understand the effect of random edge-dropping in other practical settings, e.g.
-link prediction, and even graph-level tasks.
-In general, there is a need for a broader investigation into methods designed for training deep GNNs.
-Specifically, analyzing various strategies designed for mitigating over-smoothing (see Rusch et al.
-(2023) for a review), particularly in the context of over-squashing, could be invaluable for designing deep GNNs for long-range tasks.
+Currently, real-world datasets are classified as short- or long-range tasks based on extensive model training [4] or weak proxy measures like node homophily [79].
+Developing a reliable measure of information mixing in the ground-truth data could greatly benefit the research community.
+Such a measure would enable more precise identification of short-, intermediate- and long-range tasks, improving evaluation and benchmarking.
+Another interesting direction is to investigate the significance of over-squashing in graph-level tasks, where the aggregation module of MPNNs enables some mixing of information from distant nodes.
+To the best of our knowledge, Giovanni et al.
+[33] is the only work that directly addresses this question, offering strong theoretical insights.
+However, empirical validation of these effects remains limited.
+Appendix
+[…]
 Appendix A Related Works
@@
 A popular choice for reducing over-smoothing in GNNs is to regularize the model.
-Recall that DropEdge (Rong et al., 2020) implicitly regularizes the model by adding noise to the learning trajectory (Section 2.2).
-Similarly, Dropout Srivastava et al.
-(2014) drops entries from the input features to each layer, and DropMessage (Fang et al., 2023) drops entries from the message matrix (before the aggregation step).
-Graph Drop Connect (GDC) (Hasanzadeh et al., 2020) combines DropEdge and DropMessage together, resulting in a layer-wise sampling scheme that uses a different subgraph for message-aggregation over each feature dimension.
+Recall that DropEdge [68] implicitly regularizes the model by adding noise to it (Section 2.2).
+A similarly regularization effect is observed with the methods discussed in the main text – DropNode [23], DropAgg [43], DropGNN [65], Dropout [77] and DropMessage [21].
+Graph Drop Connect (GDC) [37] combines DropEdge and DropMessage together, resulting in a layer-wise sampling scheme that uses a different subgraph for message-aggregation over each feature dimension.
 These methods successfully addressed the over-smoothing problem, enabling the training of deep GNNs, and performed competitively on several benchmarking datasets.
-Another powerful form of implicit regularization is feature normalization, which has proven crucial in enhancing the performance and stability of several types of neural networks (Huang et al., 2020).
-Exploiting the inductive bias in graph-structured data, normalization techniques like PairNorm (Zhao & Akoglu, 2020), Differentiable Group Normalization (DGN) (Zhou et al., 2020) and NodeNorm (Zhou et al., 2021b) have been proposed to reduce over-smoothing in GNNs.
-On the other hand, Energetic Graph Neural Networks (EGNNs) Zhou et al.
-(2021a) explicitly regularize the optimization by constraining the layer-wise Dirichlet energy to a predefined range.
-In a different vein, motivated by the success of residual networks (ResNets) (He et al., 2016) in computer vision, Li et al.
-(2019) proposed the use of residual connections to prevent the smoothing of representations.
+Another powerful form of implicit regularization is feature normalization, which has proven crucial in enhancing the performance and stability of several types of neural networks [41].
+Exploiting the inductive bias in graph-structured data, normalization techniques like PairNorm [93], Differentiable Group Normalization (DGN) [95] and NodeNorm [97] have been proposed to reduce over-smoothing in GNNs.
+On the other hand, Energetic Graph Neural Networks (EGNNs) [96] explicitly regularize the optimization by constraining the layer-wise Dirichlet energy to a predefined range.
+In a different vein, motivated by the success of residual networks (ResNets) [38] in computer vision, Li et al.
+[49] proposed the use of residual connections to prevent the smoothing of representations.
 Residual connections successfully improved the performance of GCN on a range of graph-learning tasks.
 Chen et al.
-(2020) introduced GCNII, which uses skip connections from the input to all hidden layers.
+[13] introduced GCN-II, which uses skip connections from the input to all hidden layers.
 This layer wise propagation rule has allowed for training of ultra-deep networks – up to 64 layers.
-Some other architectures, like the Jumping Knowledge Network (JKNet) Xu et al.
-(2018) and the Deep Adaptive GNN (DAGNN) Liu et al.
-(2020), aggregate the representations from all layers, $\{\bm{z}^{\left(\ell\right)}_{i}\}_{\ell=1}^{L}$, before processing them through a readout layer.
-A.2 Methods for Alleviating Over-squashing
+Some other architectures, like the Jumping Knowledge Network (JKNet) [84] and the Deep Adaptive GNN (DAGNN) [53], aggregate the representations from all layers, $\{\bm{z}^{\left(\ell\right)}_{i}\}_{\ell=1}^{L}$, before processing them through a readout layer.
+A.2 Homophily Bias in Evaluation of Techniques for Deep GNN
+Table 3: Statistics of node-classification datasets. Homophily measures as defined in Lim et al. [52].
+Dataset Nodes Edges Features Classes Homophily
+Homophilic Networks
+Reddit 232,965 114,615,892 602 41 0.653
+OGBN-ArXiv 169,343 1,166,243 128 40 0.416
+Coauthor-CS 18,333 163,788 6,805 15 0.755
+Coauthor-Physics 34,493 495,924 8,415 5 0.847
+Wiki-CS 11,701 216,123 300 10 0.568
+Amazon-Computers 13,752 491,722 767 10 0.700
+Amazon-Photo 7,650 238,162 745 8 0.772
+Heterophilic Networks
+Flickr 89,250 899,756 500 7 0.070
+Cornell 183 298 1,703 5 0.031
+Texas 183 325 1,703 5 0.001
+Wisconsin 251 515 1,703 5 0.094
+We examine the evaluation protocols commonly used for assessing methods aimed at alleviating over-smoothing in deep GNNs – many of which are also widely adopted for training deep architectures.
+Notably, we highlight a misalignment between the intended goal of these methods – to improve the trainability of deep GNNs – and their evaluation, which is often restricted to short-range tasks.
+For example, DropEdge [68] was evaluated on Cora [57], CiteSeer [31], PubMed [62], and a version of Reddit [36] distinct from the one used in our experiments.
+The first three exhibit high label homophily (see Table 4) and are known to be better modelled by shallower networks [95].
+Reddit also displays strong homophily, as can be seen in Table 3.
+Similarly, DropNode [23] was evaluated on Cora, CiteSeer, and PubMed; DropAgg [43] on Cora ML, CiteSeer, and OGBN-ArXiv [40], which has moderate homophily; DropMessage [68] was evaluated on Cora, CiteSeer, PubMed, OGBN-ArXiv, and Flickr, with only the latter having low homophily; GDC [37] was evaluated on Cora, Cora ML and CiteSeer.
+A similar trend can be observed in the evaluation of feature normalization techniques used to regularize GNNs.
+PairNorm [93] and DGN [95] were evaluated on Cora, CiteSeer, PubMed, and Coauthor-CS [75]; NodeNorm [97] on Cora, CiteSeer, PubMed, Coauthor-CS, Wiki-CS [58], and Amazon-Photo [75]; and EGNNs [96] on Cora, PubMed, Coauthor-Physics [75], and OGBN-ArXiv – all of these datasets are highly homophilic.
+A similar trend is observed in the evaluation of architectural modifications designed to enable deeper GNNs.
+GCN-II [13] on Cora, CiteSeer, PubMed, and Chameleon; JKNet [84] on Cora, CiteSeer, and Reddit; and DAGNN [53] on Cora, CiteSeer, PubMed, Coauthor-CS, Coauthor-Physics, Amazon-Computers [75], and Amazon-Photo – many of these datasets are highly homophilic as well.
+This pattern indicates that an overwhelming proportion of evaluations have been restricted to short-range, homophilic tasks.
+Such a narrow focus risks overstating the general effectiveness of these methods and masking their potential limitations in long-range scenarios.
+A few exceptions stand out.
+DropGNN [65], which was evaluated on graph-classification from the TUDataset [61], aligning more closely with evaluations of rewiring methods targeting over-squashing [44, 8].
+NodeNorm, while primarily evaluated on homophilic datasets, was also tested on three heterophilic graphs: Cornell, Texas, and Wisconsin [66].
+GCN-II saw broader evaluation, including on several long-range tasks such as Chameleon, Cornell, Texas, Wisconsin, and the Protein-Protein Interaction (PPI) networks [36].
+Lastly, JKNet was also evaluated on the PPI networks.
+A.3 Methods for Alleviating Over-squashing
 In this section, we will review some of the graph rewiring methods proposed to address the problem of over-squashing.
@@
 That is the graph rewiring is guided by the objective of optimizing some chosen topological properties.
-For instance, Alon & Yahav (2021) introduced a fully-adjacent (FA) layer, wherein messages are passed between all nodes.
+For instance, Alon & Yahav [4] introduced a fully-adjacent (FA) layer, wherein messages are passed between all nodes.
 GNNs using a FA layer in the final message-passing step were shown to outperform the baselines on a variety of long-range tasks, revealing the importance of information exchange between far-off nodes which standard message-passing cannot facilitate.
 Topping et al.
-(2022) proposed a curvature-based rewiring strategy, called the Stochastic Discrete Ricci Flow (SDRF), which aims to reduce the “bottleneckedness” of a graph by adding suitable edges, while simultaneously removing edges in an effort to preserve the statistical properties of the original topology.
+[79] proposed a curvature-based rewiring strategy, called the Stochastic Discrete Ricci Flow (SDRF), which aims to reduce the “bottleneckedness” of a graph by adding suitable edges, while simultaneously removing edges in an effort to preserve the statistical properties of the original topology.
 Black et al.
-(2023) proposed the Greedy Total Resistance (GTR) technique, which optimizes the graph’s total resistance by greedily adding edges to achieve the greatest improvement.
+[8] proposed the Greedy Total Resistance (GTR) technique, which optimizes the graph’s total resistance by greedily adding edges to achieve the greatest improvement.
 One concern with graph rewiring methods is that unmoderated densification of the graph, e.g.
@@
 Accordingly, Gutteridge et al.
-(2023) propose a Dynamically Rewired (DRew) message-passing framework that gradually densifies the graph.
+[35] propose a Dynamically Rewired (DRew) message-passing framework that gradually densifies the graph.
 Specifically, in a given layer $\ell$, node $i$ aggregates messages from its entire $\ell$-hop receptive field instead of just the immediate neighbors.
@@
 Some other rewiring techniques don’t exactly classify as spatial or spectral methods.
-For instance, Probabilistically Rewired MPNN (PR-MPNN) Qian et al.
-(2024) learns to probabilistically rewire a graph, effectively mitigating under-reaching as well as over-squashing.
+For instance, Probabilistically Rewired MPNN (PR-MPNN) [67] learns to probabilistically rewire a graph, effectively mitigating under-reaching as well as over-squashing.
 Finally, Gabrielsson et al.
-(2023) proposed connecting all nodes at most $r$-hops away, for some $r\in\mathbb{N}$, and introducing positional embeddings to allow for distance-aware aggregation of messages.
-A.3 Towards a Unified Treatment
+[25] proposed connecting all nodes at most $r$-hops away, for some $r\in\mathbb{N}$, and introducing positional embeddings to allow for distance-aware aggregation of messages.
+A.4 Towards a Unified Treatment
 Several studies have shown that an inevitable trade-off exists between the problems of over-smoothing and over-squashing, meaning that optimizing for one will compromise the other.
 For instance, Topping et al.
-(2022); Nguyen et al.
-(2023) showed that negatively curved edges create bottlenecks in the graph resulting in over-squashing of information.
+[79], Nguyen et al.
+[63] showed that negatively curved edges create bottlenecks in the graph resulting in over-squashing of information.
 On the other hand, Nguyen et al.
-(2023, Proposition 4.3) showed that positively curved edges in a graph contribute towards the over-smoothing problem.
+[63, Proposition 4.3] showed that positively curved edges in a graph contribute towards the over-smoothing problem.
 To address this trade-off, they proposed Batch Ollivier-Ricci Flow (BORF), which adds new edges adjacent to the negatively curved ones, and simultaneously removes positively curved ones.
 In a similar vein, Giraldo et al.
-(2023) demonstrated that the minimum number of message-passing steps required to reach a given level of over-smoothing is inversely related to the Cheeger’s constant, $h_{\mathcal{G}}$.
+[34] demonstrated that the minimum number of message-passing steps required to reach a given level of over-smoothing is inversely related to the Cheeger’s constant, $h_{\mathcal{G}}$.
 This again implies an inverse relationship between over-smoothing and over-squashing.
@@
 One such work is Karhadkar et al.
-(2023), which proposed a rewiring algorithm that adds edges to the graph but does not remove any.
-The First-order Spectral Rewiring (FoSR) algorithm computes, as the name suggests, a first order approximation to the spectral gap of $\bm{L}^{\text{sym}}$, and adds edges with the aim of maximizing it.
-Since the spectral gap directly relates to Cheeger’s constant through Cheeger’s inequality (Alon & Milman, 1985; Alon, 1986; Sinclair & Jerrum, 1989), this directly decreases the over-squashing levels.
+[44], which proposed a rewiring algorithm that adds edges to the graph but does not remove any.
+The First-order Spectral Rewiring (FoSR) algorithm computes, as the name suggests, a first-order approximation to the spectral gap of the symmetric Laplacian matrix ($\bm{L}^{\text{sym}}=\bm{I}_{N}-(\bm{D}^{\dagger})^{1/2}\bm{A}(\bm{D}^{\dagger})^{1/2}$), and adds edges with the aim of maximizing it.
+Since the spectral gap directly relates to Cheeger’s constant – a measure of bottleneck-edness in the graph – through Cheeger’s inequality [2, 3, 76], this directly decreases the over-squashing levels.
 Moreover, Karhadkar et al.
-(2023, Figure 5) empirically demonstrated that addition of (up to a small number of) edges selected by FoSR can lower the Dirichlet energy of the representations.
-Taking a completely opposite approach, CurvDrop (Liu et al., 2023) adapted DropEdge to remove negatively curved edges sampled from a distribution proportional to their curvatures.
-CurvDrop directly reduces over-squashing and, as a side benefit of operating on a sparser subgraph, also mitigates over-smoothing.
+[44, Figure 5] empirically demonstrated that addition of (up to a small number of) edges selected by FoSR can lower the Dirichlet energy of the representations, suggesting the method’s potential to simultaneously tackle over-smoothing.
+Taking a somewhat opposite approach, Liu et al.
+[54] adapted DropEdge to remove negatively curved edges sampled from a distribution proportional to edge curvatures.
+Their method, called CurvDrop, directly reduces over-squashing and, as a side benefit of operating on a sparser subgraph, also mitigates over-smoothing.
 Appendix B Proofs
-B.1 Expected Propagation Matrix under DropEdge
+B.1 Expected Propagation Matrix under DropEdge-variants
 Lemma.
@@
 masks, $m_{1},\ldots,m_{d_{i}}\sim\text{Bern}\left(1-q\right)$, on incoming edges to node $i$, the total number of messages is given by
-$1+\sum_{k=1}^{d_{i}}m_{k}=1+M_{i}$
+$1+\sum_{k=1}^{d_{i}}m_{k}=1+M_{i}$ (B.1)
 where $M_{i}\sim\text{Binom}\left(d_{i},1-q\right)$. Under asymmetric normalization (see Section 2.1), the expected weight of the message along the self-loop is computed as follows:
@@
 $=\frac{1}{d_{i}}\left(1-\mathbb{E}_{\mathsf{DE}}\left[\hat{\bm{A}}^{\left(1\right)}_{ii}\right]\right)$ (B.14)
+Analysis of DropEdge-variants.
+We will similarly derive the expected propagation matrix for other random edge-dropping algorithms.
+First off, DropNode [23] samples nodes and drops corresponding columns from the aggregation matrix directly, followed by rescaling of its entries:
+$\begin{split}\mathbb{E}_{\mathsf{DN}}\left[\frac{1}{1-q}\hat{\bm{A}}\right]=\frac{1}{1-q}\times\left(1-q\right)\hat{\bm{A}}=\hat{\bm{A}}\end{split}$ (B.15)
+That is, the expected propagation matrix is the same as in a NoDrop model ($q=0$).
+Nodes sampled by DropAgg [43] don’t aggregate messages. Therefore, if $\hat{\bm{A}}=\hat{\bm{A}}^{\text{asym}}$, then the expected propagation matrix is given by
+$\begin{split}\mathbb{E}_{\mathsf{DA}}\left[\hat{\bm{A}}_{ii}\right]&=q+\frac{1-q}{d_{i}+1}=\frac{1+d_{i}q}{d_{i}+1}>\mathbb{E}_{\mathsf{DE}}\left[\hat{\bm{A}}_{ii}\right]\\ \mathbb{E}_{\mathsf{DA}}\left[\hat{\bm{A}}_{ij}\right]&=\frac{1}{d_{i}}\left(1-\mathbb{E}_{\mathsf{DA}}\left[\hat{\bm{A}}_{ii}\right]\right)<\mathbb{E}_{\mathsf{DE}}\left[\hat{\bm{A}}_{ij}\right]\end{split}$ (B.16)
+Finally, DropGNN [65] samples nodes which neither propagate nor aggregate messages. From any node’s perspective, if it is not sampled, then its aggregation weights are computed as for DropEdge:
+$\begin{split}\mathbb{E}_{\mathsf{DG}}\left[\hat{\bm{A}}_{ii}\right]&=q+\left(1-q\right)\mathbb{E}_{\mathsf{DE}}\left[\hat{\bm{A}}_{ii}\right]=q+\frac{1-q^{d_{i}+1}}{d_{i}+1}>\mathbb{E}_{\mathsf{DA}}\left[\hat{\bm{A}}_{ii}\right]\\ \mathbb{E}_{\mathsf{DG}}\left[\hat{\bm{A}}_{ij}\right]&=\frac{1}{d_{i}}\left(1-\mathbb{E}_{\mathsf{DG}}\left[\hat{\bm{A}}_{ii}\right]\right)<\mathbb{E}_{\mathsf{DA}}\left[\hat{\bm{A}}_{ij}\right]\end{split}$ (B.17)
 B.2 Sensitivity in L-layer Linear GCNs
@@
 $\mathbb{E}_{\ldots}\left[\left(\hat{\bm{A}}^{L}\right)_{ij}\right]=\sum_{\left(u_{0},\ldots,u_{L}\right)\in\mathsf{Paths}\left(j\to i\right)}\prod_{\ell=1}^{L}\mathbb{E}_{\ldots}\left[\hat{\bm{A}}_{u_{\ell}u_{\ell-1}}\right]<\mathbb{E}_{\mathsf{ND}}\left[\left(\hat{\bm{A}}^{L}\right)_{ij}\right]$ (B.18)
-where $\mathsf{ND}$ refers to a NoDrop model ($q=0$), the placeholder $\cdots$ can be replaced with one of the edge-dropping methods $\mathsf{DE}$, $\mathsf{DA}$ or $\mathsf{DG}$, and the corresponding entries of $\mathbb{E}_{\ldots}[\hat{\bm{A}}]$ can be plugged in from Equation 3.3, Equation 3.5 and Equation 3.6, respectively.
+where $\mathsf{ND}$ refers to a NoDrop model ($q=0$), the placeholder $\cdots$ can be replaced with one of the edge-dropping methods $\mathsf{DE}$, $\mathsf{DA}$ or $\mathsf{DG}$, and the corresponding entries of $\mathbb{E}_{\ldots}[\hat{\bm{A}}]$ can be plugged in from Equation 3.3, Equation B.16 and Equation B.17, respectively.
 Moreover, the sensitivity monotonically decreases as the dropping probability is increased.
@@
 More formally, denote the set of paths connecting the two nodes by
-$\mathsf{Paths}\left(j\to i\right)=\left\{\left(u_{0},\ldots,u_{L}\right):u_{0}=j;u_{L}=i;\left(u_{\ell-1}\to u_{\ell}\right)\in\mathcal{E},\forall\ell\in\left[L\right]\right\}$
+$\mathsf{Paths}\left(j\to i\right)=\left\{\left(u_{0},\ldots,u_{L}\right):u_{0}=j;u_{L}=i;\left(u_{\ell-1}\to u_{\ell}\right)\in\mathcal{E},\forall\ell\in\left[L\right]\right\}$ (B.19)
 The $\left(i,j\right)$-entry in the propagation matrix is given by
@@
 The right hand side of the inequality is the $\left(i,j\right)$-entry in the Lth power of the propagation matrix of a NoDrop model.
-From Equation 3.5 and Equation 3.6, we know that Equation B.17 is true for DropAgg and DropGNN as well.
+From Equation B.16 and Equation B.17, we know that Equation B.22 is true for DropAgg and DropGNN as well.
 We conclude the first part of the proof using Equation 3.2 – the sensitivity of node $i$ to node $j$ is proportional to $(\dot{\bm{P}}^{L})_{ij}$.
@@
 Each of the terms on the right are increasing in $q$, hence, all the $\dot{\bm{P}}_{u_{\ell}u_{\ell-1}}$ factors are decreasing in $q$.
-Using this result with Equation B.15, we conclude the second part of the theorem.
-Appendix C Test-time Monte-Carlo Averaging
-In Section 3, we focused on the expected sensitivity of the stochastic representations in models using DropEdge-like regularization strategies.
+Similarly, $\mathbb{E}_{\mathsf{DA}}[\hat{\bm{A}}_{ij}]$ and $\mathbb{E}_{\mathsf{DG}}[\hat{\bm{A}}_{ij}]$ decrease with increasing $q$.
+Using these results with Equation B.20, we conclude the second part of the theorem.
+Appendix C Theoretical Extensions
+C.1 Sensitivity in Nonlinear MPNNs
+While linear networks are useful in simplifying the theoretical analysis, they are often not practical.
+In this subsection, we will consider the upper bounds on sensitivity established in previous works, and extend them to the DropEdge setting.
+ReLU GCNs.
+Xu et al.
+[84] considered the case of ReLU nonlinearity, so that the update rule is $\bm{Z}^{\left(\ell\right)}=\mathsf{ReLU}(\hat{\bm{A}}\bm{Z}^{\left(\ell-1\right)}\bm{W}^{\left(\ell\right)})$.
+Additionally, it makes the simplifying assumption that each path in the computational graph is active with a fixed probability, $\rho$ [45, Assumption A1p-m].
+Accordingly, the sensitivity (in expectation) between any two nodes is given as
+$\left\|\mathbb{E}_{\mathsf{ReLU}}\left[\frac{\partial\bm{z}^{\left(L\right)}_{i}}{\partial\bm{x}_{j}}\right]\right\|_{1}=\left[\rho\left\|\prod_{\ell=1}^{L}\bm{W}^{\left(\ell\right)}\right\|_{1}\right]\left(\hat{\bm{A}}^{L}\right)_{ij}=\zeta_{1}^{\left(L\right)}\left(\hat{\bm{A}}^{L}\right)_{ij}$ (C.1)
+where $\zeta_{1}^{\left(L\right)}$ depends only on the depth $L$, and is independent of the choice of nodes $i,j\in\mathcal{V}$. Taking an expectation w.r.t. the random edge masks, we get
+$\mathbb{E}_{\mathbf{M}^{\left(1\right)},\ldots,\mathbf{M}^{\left(L\right)}}\left[\left\|\mathbb{E}_{\mathsf{ReLU}}\left[\frac{\partial\bm{z}^{\left(L\right)}_{i}}{\partial\bm{x}_{j}}\right]\right\|_{1}\right]$ $=\zeta_{1}^{\left(L\right)}\left(\mathbb{E}\left[\prod_{\ell=1}^{L}\hat{\bm{A}}^{\left(\ell\right)}\right]\right)_{ij}=\zeta_{1}^{\left(L\right)}\left(\mathbb{E}\left[\hat{\bm{A}}\right]^{L}\right)_{ij}$ (C.2)
+Using Theorem 3.1, we conclude that in a ReLU-GCN, DropEdge, DropAgg and DropGNN will reduce the expected sensitivity between nodes L-hops away.
+Empirical observations in s 1(a) and 4 suggest that we may expect an increase in sensitivity to neighboring nodes, but a significant decrease in sensitivity to those farther away.
+Source-only Message Functions. Black et al. [8, Lemma 3.2] considers MPNNs with aggregation functions of the form
+$\mathsf{Agg}^{\left(\ell\right)}\left(\bm{z}_{i}^{\left(\ell-1\right)},\left\{\bm{z}_{j}^{\left(\ell-1\right)}:j\in\mathbb{S}^{\left(1\right)}\left(i\right)\right\}\right)=\sum_{j\in\mathbb{B}^{\left(1\right)}\left(i\right)}\hat{\bm{A}}_{ij}\mathsf{Msg}^{\left(\ell\right)}\left(\bm{z}_{j}^{\left(\ell-1\right)}\right)$ (C.3)
+and $\mathsf{Upd}$ and $\mathsf{Msg}$ functions with bounded gradients. In this case, the sensitivity between two nodes $i,j\in\mathcal{V}$ can be bounded as
+$\left\|\frac{\partial\bm{z}^{\left(L\right)}_{i}}{\partial\bm{x}_{j}}\right\|_{1}\leq\zeta_{2}^{\left(L\right)}\left(\sum_{\ell=0}^{L}\hat{\bm{A}}^{\ell}\right)_{ij}$ (C.4)
+As before, we can use the independence of edge masks to get an upper bound on the expected sensitivity:
+$\mathbb{E}_{\mathbf{M}^{\left(1\right)},\ldots,\mathbf{M}^{\left(L\right)}}\left[\left\|\frac{\partial\bm{z}^{\left(L\right)}_{i}}{\partial\bm{x}_{j}}\right\|_{1}\right]$ $\leq\zeta_{2}^{\left(L\right)}\left(\mathbb{E}\left[I_{N}+\sum_{\ell=1}^{L}\prod_{k=1}^{\ell}\hat{\bm{A}}^{\left(k\right)}\right]\right)_{ij}=\zeta_{2}^{\left(L\right)}\left(\sum_{\ell=0}^{L}\mathbb{E}\left[\hat{\bm{A}}\right]^{\ell}\right)_{ij}$ (C.5)
+Figure 5 shows the plot of the entries of $\sum_{\ell=0}^{6}\dot{\bm{P}}^{\ell}$ (i.e.
+for DropEdge), as in the upper bound above, with $\hat{\bm{A}}=\hat{\bm{A}}^{\text{asym}}$.
+We observe that the sensitivity between nearby nodes marginally increases, while that between distant nodes notably decreases (similar to 1(a)), suggesting significant over-squashing.
+Similar observations can be made with $\hat{\bm{A}}=\hat{\bm{A}}^{\text{sym}}$, and for other DropEdge-variants.
+Source-and-Target Message Functions. Topping et al. [79, Lemma 1] showed that if the aggregation function is instead given by
+$\mathsf{Agg}^{\left(\ell\right)}\left(\bm{z}_{i}^{\left(\ell-1\right)},\left\{\bm{z}_{j}^{\left(\ell-1\right)}:j\in\mathbb{S}^{\left(1\right)}\left(i\right)\right\}\right)=\sum_{j\in\mathbb{B}^{\left(1\right)}\left(i\right)}\hat{\bm{A}}_{ij}\mathsf{Msg}^{\left(\ell\right)}\left(\bm{z}_{i}^{\left(\ell-1\right)},\bm{z}_{j}^{\left(\ell-1\right)}\right)$ (C.6)
+then the sensitivity between nodes $i\in\mathcal{V}$ and $j\in\mathbb{S}^{\left(L\right)}\left(i\right)$ can be bounded as
+$\left\|\frac{\partial\bm{z}^{\left(L\right)}_{i}}{\partial\bm{x}_{j}}\right\|_{1}\leq\zeta_{3}^{\left(L\right)}\left(\hat{\bm{A}}^{L}\right)_{ij}$ (C.7)
+With random edge-dropping, this bound can be adapted as follows:
+$\mathbb{E}_{\mathbf{M}^{\left(1\right)},\ldots,\mathbf{M}^{\left(L\right)}}\left[\left\|\frac{\partial\bm{z}^{\left(L\right)}_{i}}{\partial\bm{x}_{j}}\right\|_{1}\right]\leq\zeta_{3}^{\left(L\right)}\left(\mathbb{E}\left[\hat{\bm{A}}\right]^{L}\right)_{ij}$ (C.8)
+which is similar to Equation C.2, only with a different proportionality constant, that is anyway independent of the choice of nodes.
+Here, again, we invoke Theorem 3.1 to conclude that $(\mathbb{E}[\hat{\bm{A}}]^{L})_{ij}$ decreases monotonically with increasing DropEdge probability $q$.
+This implies that, in a non-linear MPNN with $\hat{\bm{A}}=\hat{\bm{A}}^{\text{asym}}$, DropEdge lowers the sensitivity bound given above.
+Empirical results in Figure 4 support the same conclusion for $\hat{\bm{A}}=\hat{\bm{A}}^{\text{sym}}$.
+C.2 Test-time Monte-Carlo Dropout
+Up until now, we have focused on the expected sensitivity of the stochastic representations in models using DropEdge-variants.
 This corresponds to their training-time behavior, wherein the activations are random.
 At test-time, the standard practice is to turn these methods off by setting $q=0$.
-However, this raises the over-smoothing levels back up Xuanyuan et al.
-(2023).
+However, this raises the over-smoothing levels back up [86].
 Another way of making predictions is to perform multiple stochastic forward passes, as during training, and then averaging the model outputs.
-This is similar to Monte-Carlo Dropout, which is an efficient way of ensemble averaging in MLPs Gal & Ghahramani (2016b), CNNs Gal & Ghahramani (2016a) and RNNs Gal & Ghahramani (2016c).
-In addition to alleviating over-smoothing, this approach also outperforms the standard implementation Xuanyuan et al.
-(2023).
+This is similar to Monte-Carlo Dropout, which is an efficient way of ensemble averaging in MLPs [27], CNNs [26] and RNNs [28].
+In addition to alleviating over-smoothing, this approach also outperforms the standard implementation in practical settings [86].
 We can study the effect of random edge-dropping in this setting by examining the sensitivity of the expected representations:
-$\left\|\frac{\partial}{\partial\bm{x}_{j}}\mathbb{E}\left[\bm{z}^{\left(L\right)}_{i}\right]\right\|_{1}$
-In linear models, the order of the two operations – expectation and sensitivity computation – is irrelevant:
-$\mathbb{E}\left[\left\|\frac{\partial\bm{z}^{\left(L\right)}_{i}}{\partial\bm{x}_{j}}\right\|_{1}\right]=\left\|\mathbb{E}\left[\hat{\bm{A}}_{ij}\right]\bm{W}\right\|_{1}=\left\|\mathbb{E}\left[\frac{\partial\bm{z}^{\left(L\right)}_{i}}{\partial\bm{x}_{j}}\right]\right\|_{1}=\left\|\frac{\partial}{\partial\bm{x}_{j}}\mathbb{E}\left[\bm{z}^{\left(L\right)}_{i}\right]\right\|_{1}$ (C.1)
+$\left\|\frac{\partial}{\partial\bm{x}_{j}}\mathbb{E}\left[\bm{z}^{\left(L\right)}_{i}\right]\right\|_{1}=\left\|\mathbb{E}\left[\frac{\partial\bm{z}^{\left(L\right)}_{i}}{\partial\bm{x}_{j}}\right]\right\|_{1}$ (C.9)
+In linear models, the order of the two operations – expectation and 1-norm – is irrelevant:
+$\left\|\mathbb{E}\left[\frac{\partial\bm{z}^{\left(L\right)}_{i}}{\partial\bm{x}_{j}}\right]\right\|_{1}=\left\|\mathbb{E}\left[\left(\hat{\bm{A}}^{L}\right)_{ij}\right]\bm{W}\right\|_{1}=\mathbb{E}\left[\left(\hat{\bm{A}}^{L}\right)_{ij}\left\|\bm{W}\right\|_{1}\right]=\mathbb{E}\left[\left\|\frac{\partial\bm{z}^{\left(L\right)}_{i}}{\partial\bm{x}_{j}}\right\|_{1}\right]$ (C.10)
 In general, the two quantities can be related using the convexity of norms and Jensen’s inequality:
-$\left\|\frac{\partial}{\partial\bm{x}_{j}}\mathbb{E}\left[\bm{z}^{\left(L\right)}_{i}\right]\right\|_{1}\leq\mathbb{E}\left[\left\|\frac{\partial\bm{z}^{\left(L\right)}_{i}}{\partial\bm{x}_{j}}\right\|_{1}\right]$ (C.2)
-Therefore, the discussion in the previous subsections extends to the MC-averaged representations as well. Although tighter bounds may be derived for this setting, we will leave that for future works.
-Appendix D Additional Figures
-In this section, we present some additional figures that demonstrate the negative effects of random edge-dropping, particularly focusing on providing empirical evidence for scenarios not covered by the theory in Section 3.
-Additionally, we provide visual evidence of the negative effects of DropNode, despite the fact that it preserves sensitivity between nodes (in expectation).
+$\left\|\frac{\partial}{\partial\bm{x}_{j}}\mathbb{E}\left[\bm{z}^{\left(L\right)}_{i}\right]\right\|_{1}\leq\mathbb{E}\left[\left\|\frac{\partial\bm{z}^{\left(L\right)}_{i}}{\partial\bm{x}_{j}}\right\|_{1}\right]\leq\ldots$ (C.11)
+Therefore, the upper bound results in Section C.1 trivially extend to the MC-averaged representations. Although tighter bounds may be derived for this setting, we leave that for future works.
+Appendix D Empirical Sensitivity Analysis
+In this section, we present some supplemental figures demonstrating the negative effects of random edge-dropping, particularly focusing on scenarios not covered by the theory.
+We also elaborate on the setup used for the empirical sensitivity analysis in Section 3.
 D.1 Symmetrically Normalized Propagation Matrix
 Figure 4: Entries of $\ddot{\bm{P}}^{6}$, averaged after binning node-pairs by their shortest distance.
-The results in Section 3.1 correspond to the use of $\hat{\bm{A}}=\hat{\bm{A}}^{\text{asym}}$ for aggregating messages – in each message passing step, only the in-degree of node $i$ is used to compute the aggregation weights of the incoming messages.
-In practice, however, it is more common to use the symmetrically normalized propagation matrix, $\hat{\bm{A}}=\hat{\bm{A}}^{\text{sym}}$, which ensures that nodes with high degree do not dominate the information flow in the graph Kipf & Welling (2017).
+The results in Section 3 correspond to the use of $\hat{\bm{A}}=\hat{\bm{A}}^{\text{asym}}$ for aggregating messages – in each message passing step, only the in-degree of node $i$ is used to compute the aggregation weights of the incoming messages.
+In practice, however, it is more common to use the symmetrically normalized propagation matrix, $\hat{\bm{A}}=\hat{\bm{A}}^{\text{sym}}$, which ensures that nodes with high out-degree do not dominate the information flow in the graph [47].
 As in Equation 3.2, we are looking for
-$\mathbb{E}_{\mathbf{M}^{\left(1\right)},\ldots,\mathbf{M}^{\left(L\right)}}\left[\prod_{l=1}^{L}\hat{\bm{A}}^{\left(\ell\right)}\right]=\ddot{\bm{P}}^{L}$
+$\ddot{\bm{P}}^{L}\coloneqq\mathbb{E}_{\mathbf{M}^{\left(1\right)},\ldots,\mathbf{M}^{\left(L\right)}}\left[\prod_{\ell=1}^{L}\hat{\bm{A}}^{\left(\ell\right)}\right]$ (D.1)
 where $\ddot{\bm{P}}\coloneqq\mathbb{E}_{\mathsf{DE}}[\hat{\bm{A}}^{\text{sym}}]$.
 While $\ddot{\bm{P}}$ is analytically intractable, we can approximate it using Monte-Carlo sampling.
-Accordingly, we use 20 samples of $\mathbf{M}$ to compute an approximation of $\ddot{\bm{P}}$, and plot out the entries of $\ddot{\bm{P}}^{L}$, as we did for $\dot{\bm{P}}^{L}$ in Figure 1.
-The results are presented in Figure 4, which shows that while the sensitivity between nearby nodes is affected to a lesser extent compared to those observed in Figure 1, that between far-off nodes is significantly reduced, same as earlier.
+Accordingly, we use the Cora dataset, and sample 20 DropEdge masks to compute an approximation of $\ddot{\bm{P}}$, and plot out the entries of $\ddot{\bm{P}}^{L}$, as we did for $\dot{\bm{P}}^{L}$ in 1(a).
+The results are presented in Figure 4, which shows that while the sensitivity between nodes up to 3 hops away is increased, that between nodes farther off is significantly reduced, same as in 1(a).
 D.2 Upper Bound on Expected Sensitivity
-Figure 5: Entries of $\sum_{l=0}^{6}\dot{\bm{P}}^{l}$, averaged after binning node-pairs by their shortest distance.
+Figure 5: Entries of $\sum_{\ell=0}^{6}\dot{\bm{P}}^{\ell}$, averaged after binning node-pairs by their shortest distance.
 Black et al.
-(2023) showed that the sensitivity between any two nodes in a graph can be bounded using the sum of the powers of the propagation matrix.
-In Section 3.2, we extended this bound to random edge-dropping methods with independent edge masks smapled in each layer:
-$\mathbb{E}_{\mathbf{M}^{\left(1\right)},\ldots,\mathbf{M}^{\left(L\right)}}\left[\left\|\frac{\partial\bm{z}^{\left(L\right)}_{i}}{\partial\bm{x}_{j}}\right\|_{1}\right]=\zeta_{3}^{\left(L\right)}\left(\sum_{\ell=0}^{L}\mathbb{E}\left[\hat{\bm{A}}\right]^{\ell}\right)_{ij}$
-Although this bound does not have a closed form, we can use real-world graphs to study its entries.
-We randomly sample 100 molecular graphs from the Proteins dataset Dobson & Doig (2003) and plot the entries of $\sum_{l=0}^{6}\dot{\bm{P}}^{\ell}$ (corresponding to DropEdge) against the shortest distance between node-pairs.
+[8] showed that the sensitivity between any two nodes in a graph can be bounded using the sum of the powers of the propagation matrix.
+In Section C.1, we extended this bound to random edge-dropping methods with independent edge masks sampled in each layer:
+$\mathbb{E}_{\mathbf{M}^{\left(1\right)},\ldots,\mathbf{M}^{\left(L\right)}}\left[\left\|\frac{\partial\bm{z}^{\left(L\right)}_{i}}{\partial\bm{x}_{j}}\right\|_{1}\right]\leq\zeta_{3}^{\left(L\right)}\left(\sum_{\ell=0}^{L}\mathbb{E}\left[\hat{\bm{A}}\right]^{\ell}\right)_{ij}$
+Although this bound does not have a closed form, we can again use the Cora network to study its entries.
+We plot the entries of $\sum_{\ell=0}^{6}\dot{\bm{P}}^{\ell}$, corresponding to DropEdge, against the shortest distance between node-pairs.
 The results are presented in Figure 5.
-We observe a polynomial decline in sensitivity as the DropEdge probability increases, suggesting that it is unsuitable for capturing LRIs.
-D.3 Test Accuracy versus DropNode Probability
-(a) Homophilic datasets
-(b) Heterophilic datasets
+We observe an exponential decrease in the sensitivity bound as the distance between nodes increases, suggesting that DropEdge is not suitable for capturing LRIs.
+D.3 MC-Approximation of Sensitivity in Nonlinear MPNNs
+Given a target node from the Cora dataset [57], we computed the sensitivity of its representation to source nodes up to $L=6$ hops away in ReLU-GCNs of width 32.
+The raw sensitivities were normalized to obtain influence scores [84].
+This was repeated for 25 target nodes, and 25 model--dropout samples were used for each of them.
+The source nodes were binned by the shortest distance from the corresponding target node, and the influence scores were averaged over each bin to obtain an average influence from nodes $\ell$-hops away.
+Why influence scores?
+Appendix E Experiments Details
+In this section, we expand on the details of the experiments in Section 5.
+All experiments were run on a server equipped with an Intel(R) Xeon(R) E5-2620 v3 CPU, 62 GB of RAM, 4 $\times$ NVIDIA GeForce GTX TITAN X GPU (12 GB VRAM each), and CUDA version 12.4.
+E.1 Descriptions of the Datasets
+[…]
+Table 4: Statistics of node-classification datasets. Homophily measures from Lim et al. [52].
+Dataset Nodes Edges Features Classes Homophily
+Homophilic Networks
+Cora 2,708 10,556 1,433 7 0.766
+CiteSeer 3,327 9,104 3,703 6 0.627
+PubMed 19,717 88,648 500 3 0.664
+Heterophilic Networks
+Chameleon 2,277 36,051 2,325 5 0.062
+Squirrel 5,201 216,933 2,089 5 0.025
+Actor 7,600 29,926 931 5 0.011
+TwitchDE 9,498 306,276 128 2 0.142
+Node-classification Tasks.
+Cora [57], CiteSeer [31] and PubMed [62] are citation networks – their nodes represent scientific publications and an edge between two nodes indicates that one of them has cited the other.
+The features of each publication are represented by a binary vector, where each index indicates whether a specific word from a dictionary is present or absent.
+Several studies have showed that these datasets have high homophily in node labels [52, 98] and that they are modelled much better by shallower networks than by deeper ones [95].
+Chameleon and Squirrel [69] are networks of English Wikipedia web pages on the respective topics, and the edges between web pages indicate links between them.
+The task is to predict the average-monthly traffic on each of the web pages.
+The Actor dataset is induced from a larger film-director-actor-writer network [66].
+It is a network of film actors, with edges between those that occur on the same Wikipedia page, and node features are binary vectors denoting the presence of specific keywords in the corresponding Wikipedia entries.
+The task is to classify actors into five categories based on the content of their Wikipedia pages.
+Finally, TwitchDE [69] is a network of Twitch users in Germany, with the edges between them representing their mutual follower relationships.
+The node features are embeddings of the games played by the users, and the task is to predict whether the users use explicit language.
+Table 5: Statistics of graph-classification datasets.
+Dataset Graphs Nodes Edge Features Classes
+Mutag 188  17.9  39.6 7 2
+Proteins 1,113  39.1  145.6 3 2
+Enzymes 600  32.6  124.3 3 6
+Collab 5,000  74.5  4914.4 0 3
+IMDb 1,000  19.8  193.1 0 2
+Reddit 2,000  429.6  995.5 0 2
+Graph-classification Tasks.
+Following Black et al.
+[8], Karhadkar et al.
+[44], we use the graph-classification datasets introduced in Morris et al.
+[61], which were hypothesized to be long-range tasks.
+On one hand we have the molecular datasets Mutag, Proteins and Enzymes, and on another we have the social networks Reddit-Binary, IMDb-Binary and Collab.
+Their statistics are presented in Table 5.
+Mutag [17] consists of nitroaromatic compounds, with the objective of predicting their mutagenicity, i.e.
+the ability to cause genetic mutations, in cells in Salmonella typhimurium.
+Each compound is represented as a graph, where nodes correspond to atoms, represented by their type using one-hot encoding, and edges denote the bonds between them.
+Proteins [19, 9] is a collection of proteins classified as enzymes or not.
+The molecules are represented as a graph of amino acids, with edges between those separated up to 6Å apart.
+The Enzymes [9, 73] dataset is represented similarly, and the task is to classify the proteins into one of 6 Enzyme Commission (EC) numbers – a system to classify enzymes based on the reactions they catalyze.
+Collab [87] is a scientific collaboration dataset where each graph represents a researcher’s ego network.
+Nodes correspond to researchers and their collaborators, with edges indicating co-authorship.
+Each network is labeled based on the researcher’s field, which can be High Energy Physics, Condensed Matter Physics, or Astrophysics.
+IMDb-Binary [87] is a movie collaboration dataset containing the ego-networks of 1,000 actors and actresses from IMDB.
+In each graph, nodes represent actors, with edges connecting those who have co-starred in the same film.
+Each graph is derived from either the Action genre or Romance.
+Finally, Reddit-Binary [87] comprises graphs representing online discussions on Reddit, where nodes correspond to users and edges indicate interactions through comment responses.
+Each graph is labeled based on whether it originates from a Q&A or discussion-based subreddit.
+For the molecular datasets, we use the node features supplied by PyG [24], but since they are unavailable for the social networks, we set scalar features $\bm{x}_{i}=\left(1\right)$ for all nodes in these datasets, following [44].
+E.2 Training Configurations
+Dataset Splits.
+For the SyntheticZINC task, we use the train-val-test splits provided in PyG.
+For the homophilic (citation) networks, we use the ‘full’ split [12], as provided in PyG, and for the heterophilic networks, we randomly sample 60% of the nodes for training, 16% for validation, and 24% for testing.
+On the other hand, for the graph classification tasks, we sample 80% of the graphs for training, and 10% each for validation and testing, following [44, 8].
+Model Architecture.
+We standardize most of the hyperparameters across all experiments to isolate the effect of random dropping.
+Specifically, we use symmetric normalization of the adjacency matrix to compute the edge weights for GCN, and we set the number of attentions heads for GAT to 2 in order to keep the computational load manageable, while at the same time harnessing the expressiveness of the multi-headed self-attention mechanism.
+For the SyntheticZINC dataset, we fix the size of the hidden representations at 16, while we fix them to 64 for all the real-world datasets.
+In all settings, a linear transformation is applied to the node features before message-passing.
+Afterwards, a bias term is added and then the ReLU nonlinearity is applied.
+Finally, a linear readout layer is used to compute the regressand (for regression tasks) or logits (for classification).
+Dropping Probability.
+For the synthetic datasets, we experiment with a NoDrop baseline, and DropEdge, Dropout and DropMessage, each with $q=0.2$ and $q=0.5$.
+For the real-world datasets, the dropping probabilities are varied as $q=0.1,0.2,\ldots,0.9$, so as to reliably find the best performing configuration.
+We adopt the common practice of turning the dropping methods off at test-time ($q=0$), isolating the effects on optimization and generalization, which our theory does not address.
+DropSens Configurations.
+For DropSens, we use 4 possible values for proportion of information preserved over corss-edges, $c=0.5,0.8,0.9,0.95$.
+Since the dropping probability, $q_{i}$, increases with the in-degree of the target node, $d_{i}$, the proportion of all edges dropped could become very high, especially with large $c$.
+Therefore, we clip the value of $q_{i}$ by 4 possible choices, $q_{i}\leq q_{\max}\in\left\{0.2,0.3,0.5,0.8\right\}$.
+We exclude the following configurations: $\left(c,q_{\max}\right)\in\left\{\left(0.5,0.2\right),\left(0.5,0.3\right),\left(0.5,0.5\right),\left(0.8,0.2\right),\left(0.8,0.3\right)\right\}$, since they use the same dropping probability ($=q_{\max}$) for each edge, and are therefore equivalent to DropEdge.
+In summary, we test with a total of 11 configurations for DropSens to find the best one for each task.
+Optimization Algorithm.
+The models are trained using the Adam optimizer [46].
+On the SyntheticZINC dataset, the models are trained with a learning rate of $2\times 10^{-3}$ and a weight decay of $1\times 10^{-4}$, for a total of 200 epochs.
+On the real-world datasets, we use a learning rate of $1\times 10^{-3}$ and no weight decay, following Karhadkar et al.
+[44], Black et al.
+[8].
+Here, we cap the maximum number of epochs at 300.
+In both cases, the learning rate is reduced by a factor of $1\times 10^{-1}$ if the validation loss fails to improve beyond a relative threshold of $1\times 10^{-4}$ for 10 epochs, again following Karhadkar et al.
+[44], Black et al.
+[8].
+Number of Independent Runs.
+We perform only 10 independent runs on the SyntheticZINC dataset due to its consistently low variance in performance, as also observed by Giovanni et al.
+[33].
+For real-world datasets, we conduct 20 runs to identify the best-performing dropping configurations.
+We then perform a one-sided t-test to assess whether dropout improves performance, using a 90% confidence-level ($\alpha=0.1$) and targeting a statistical power of 0.9 ($\beta=0.1$).
+Under the assumption that the dropping method offers superior performance, detecting a medium effect size of $0.5$ [14] requires approximately 53 samples per group according to standard power analysis.
+We round this to 50, and accordingly perform 30 additional runs for the final comparison of the best-performing dropping configuration with the NoDrop baseline.
+E.3 DropSens Implementation
+import warnings
+import sympy
+from sympy.abc import x as q
+import torch
+from torch_geometric.utils import degree, contains_self_loops
+def drop_sens(
+edge_index: torch.Tensor,
+c: float,
+max_drop_prob: float = None
+if max_drop_prob is None:
+max_drop_prob = 1.
+# Assuming edge index does not have self loops
+if contains_self_loops(edge_index):
+warnings.warn("Degree computation in DropSens assumes absence",
+"of self-loops, but the edge_index passed has them.")
+degrees = degree(edge_index[1]).int() # Node index -> in-degree
+ds = torch.unique(degrees).tolist() # Sorted array
+mapper = torch.nan * torch.ones(ds[-1]+1)
+mapper[ds] = max_drop_prob # Node degree -> dropping prob
+for d_i in ds:
+q_i = float(sympy.N(sympy.real_roots(
+(1-c)*d_i*(1-q) - q + q**(d_i+1))[-2] # Following Equation 4.1
+)) if d_i > 0 else 0.
+if q_i > max_drop_prob:
+# Because q monontonic wrt d, and ds is sorted
+break
+mapper[d_i] = q_i
+in_degrees = degrees[edge_index[1]] # Edge index -> in-degree of target node
+qs = mapper[in_degrees] # Edge index -> dropping probability
+edge_mask = qs <= torch.rand(edge_index.size(1))
+edge_index = edge_index[:, edge_mask]
+return edge_index, edge_mask
+Listing 1: DropSens Implementation
+In 1, we present the DropSens implementation used in our experiments, relying mainly on SymPy [59].
+Figure 6: Edge-wise dropping probabilities under DropSens for varying values of $c$, along with corresponding approximations as in Equation E.1.
+Unfortunately, computing the roots of Equation 4.1 becomes slow when the in-degree $d_{i}$ is large – a common scenario in large networks.
+This issue is especially pronounced when the proportion of information preserved $c$ is large, as the dropping threshold is only met at a higher value of $d_{i}$.
+To address this computational challenge, we propose an approximation:
+$1-c=\frac{q_{i}-q_{i}^{d_{i}+1}}{d_{i}\left(1-q_{i}\right)}\approx\frac{q_{i}}{d_{i}\left(1-q_{i}\right)}\quad\Longrightarrow\quad q_{i}\approx\frac{\left(1-c\right)d_{i}}{1+\left(1-c\right)d_{i}}$ (E.1)
+This approximation becomes increasingly accurate as $c$ increases – since more information needs to be preserved, $q_{i}$ needs to be small, and hence, $q_{i}^{d_{i}+1}\to 0$.
+Figure 6 shows the DropSens probabilities for masking incoming messages based on the in-degree of the target nodes, along with the corresponding approximations.
+It is clear to see that the approximation gets increasingly more accurate for lower proportions of information loss.
+Appendix F Supplementary Experiments
+F.1 Test Accuracy versus DropEdge Probability
+Figure 7: Dropping probability versus test accuracy of DropEdge-GCN.
+The theory the explains the contrasting trends as follows: random edge-dropping pushes models to fit to local information during training, which is suitable for short-range tasks, but harms test-time performance in long-range ones.
+In Section 3, we studied the effect of edge-dropping probability on sensitivity between nodes at different distances.
+However, this analysis may be insufficient to precisely predict the impact on model performance since DropEdge-variants significantly affect the optimization trajectory as well.
+To learn more about the relationship between test-time performance and dropping probability, we evaluate DropEdge-GCNs on the heterophilic datasets; the results are shown in Figure 7.
+Clearly, on Chameleon, Squirrel and TwitchDE, the performance degrades with increasing dropping probability, as was suggested by Theorem 3.1 and 1(a).
+Surprisingly, the trends are significantly monotonic with GCNs of all depths, $L=2,4,6,8$.
+F.2 Remark on DropNode
 Figure 8: Dropping probability versus test accuracy of DropNode-GCN.
-In Equation 3.4, we noted that the expectation of sensitivity remains unchanged when using DropNode.
-However, these results were only in expectation.
-In practice, a high DropNode probability will result in poor communication between distant nodes, preventing the model from learning to effectively model LRIs.
-This is supported by the results in Table 2, where we observed a negative correlation between the test accuracy and DropNode probability.
-Moreover, DropNode was the only algorithm which recorded negative correlations on homophilic datasets.
-In Figure 6, we visualize these relationships, noting the stark contrast with Figure 2, particularly in the trends with homophilic datasets.
-Appendix E Experiment Details
-In this section, we expand on the details of the experiment in Section 4.
-Descriptions of the Datasets.
-Cora McCallum et al.
-(2000) and CiteSeer Giles et al.
-(1998) are citation networks – their nodes represent scientific publications and an edge between two nodes indicates that one of them has cited the other.
-The features of each publication are represented by a binary vector, where each index indicates whether a specific word from a dictionary is present or absent.
-Several studies have showed that these datasets have high homophily in node labels Lim et al.
-(2021); Zhu et al.
-(2020) and that they are modelled much better by shallower networks than by deeper ones Zhou et al.
-(2020).
-Chameleon and Squirrel Rozemberczki et al.
-(2021) are networks of English Wikipedia web pages on the respective topics, and the edges between web pages indicate links between them.
-The task is to predict the average-monthly traffic on each of the web pages.
-Finally, TwitchDE Rozemberczki et al.
-(2021) is a network of Twitch users in Germany, with the edges between them representing their mutual follower relationships.
-The node features are embeddings of the games played by the users.
-The task is to predict whether the users use explicit language.
-Training Hyperparameters.
-We standardise most of the hyperparameters across all experiments in order to isolate the effect of dropping probability.
-Specifically, we fix the size of the hidden representations in each layer at 64, and a linear readout layer is used to compute the node-level logits.
-The models are trained using the Adam optimizer Kingma & Ba (2015), with a learning rate of $3\times 10^{-3}$ and a weight decay of $5\times 10^{-4}$, for a total of 300 epochs.
-For GCN, we use symmetric normalization of the adjacency matrix to compute the edge weights Kipf & Welling (2017).
-GAT Runs.
-It is quite problematic to train deep GAT models due to vanishing gradients (Dasoulas et al., 2021).
-Accordingly, we discard the runs where the model fails to learn, and performs just as well as a random classifier.
-Specifically, we compute the class distribution in each of the networks, and discard the runs where the test performance does not exceed the maximum proportion.
-This comes out to be $0.3021$, $0.2107$, $0.2288$, $0.2003$ and $0.6045$ for Cora, CiteSeer, Chameleon, Squirrel and TwitchDE, respectively.
-E.1 Extension to GraphSAGE and ResGCN
-Our exploration in this work has been mainly limited to stochastic regularization methods, including edge-dropping methods, and related ones like Dropout Srivastava et al.
-(2014), DropMessage Fang et al.
-(2023) and SkipNode Lu et al.
-(2024).
-We may follow a similar intuition to study other methods as well.
-For example, GraphSAGE Hamilton et al.
-(2017) operates by sampling a subset of neighbors for message aggregation in each step.
-From an over-squashing perspective, it is very similar to DropEdge and therefore, we can expect a similar decline in sensitivity between distant nodes when using GraphSAGE.
-Figure 7: Influence distributions of a regular GCN and a ResGCN. Residual connections reduce the influence of nodes farther away by amplifying the influence of nodes close by.
-On the other hand, residual connection reduce the influence of distant nodes by increasing the influence of nearby neighbors.
-For example, in the case of ResGCN, the operation $\bm{Z}^{\left(\ell+1\right)}=\hat{\bm{A}}\bm{Z}^{\left(\ell\right)}\bm{W}^{\left(\ell+1\right)}+\bm{Z}^{\left(\ell\right)}$ first collects information from nodes $(\ell+1)$-hops away, like in a regular GCN model, and then adds information from nodes up to $\ell$-hops away, effectively decreasing the influence of the nodes $(\ell+1)$-hops away.
-In other words, we may expect residual connections to be similarly unsuited for long-range tasks, not by reducing the strength of the signal over long distances, but rather by amplifying the strength of short-range signals.
-In Figure 7, we present the empirical influence distribution Xu et al.
-(2018) of a 6-layer ResGCN, computed using the Cora network.
-Here, we can clearly see that residual connections significantly reduce the influence of distant nodes on final representations, and can thus be expected to be unsuitable for long-range tasks.
+In Equation B.15, we noted that DropNode does not suffer from loss in sensitivity.
+However, those results were in expectation.
+Moreover, our analysis did not account for the effects on the learning trajectory.
+In practice, a high DropNode probability would make it hard for information in the node features to reach distant nodes.
+This would prevent the model from learning to effectively combine information from large neighborhoods, harming generalization.
+In Figure 8, we visualize the relationship between test-time performance and DropNode probability.
+The performance monotonically decreases with increasing dropping probability, as was observed with DropEdge.
+F.3 Over-squashing or Under-fitting?
+Figure 9: DropEdge probability versus training accuracy of GCNs.
+The training performance improves with $q$, suggesting that the models are not underfitting.
+Instead, the reason for poor test-time performance (Figure 7) is that models are over-fitting to short-range signals during training, resulting in poor generalization.
+The results in the previous subsection suggest that using random edge-dropping to regularize model training leads to poor test-time performance.
+We hypothesize that this occurs because the models struggle to propagate information over long distances, causing node representations to overfit to local neighborhoods.
+However, a confounding effect is at play: DropEdge variants reduce the generalization gap by preventing overfitting to the training set, i.e.
+poorer training performance.
+If this regularization is too strong, it could lead to underfitting, which could also explain the poor test-time performance on heterophilic datasets.
+This concern is particularly relevant because the heterophilic networks are much larger than homophilic ones (see Table 4), making them more prone to underfitting.
+To investigate this, we plot the training accuracies of deep DropEdge-GCNs on the heterophilic datasets; Figure 9 shows the results.
+It is clear that the models do not underfit as the dropping probability increases.
+In fact, somewhat unexpectedly, the training metrics improve.
+Together with the results in Figure 7, we conclude that DropEdge-like methods are detrimental in long-range tasks since they cause overfitting to short-range artifacts in the training data, resulting in poor generalization at test-time.
+Appendix G Supplementary Experimental Results
+G.1 Performance of GAT with Dropping Methods
+In Table 6, we present the results of experiments in Section 5.2, but with the GAT architecture.
+For node classification tasks, we see the same dichotomy as in 1(a), with dropping methods significantly improving performance on homophilic networks, while being detrimental to performance on heterophilic networks.
+On graph classification tasks, the dropping methods improve performance, but the improvement (if any) is not statistically significant (in $28/36\approx 78\%$ cases).
+Note that the GAT architecture was unable to learn the Collab dataset, i.e.
+the performance in all cases was as good as a random classifier’s.
+Table 6: Difference in mean test accuracy (%) between the best performing configuration of each dropout method and the baseline NoDrop model, with GAT as the base model.
+Cell colors represent p-values from a t-test evaluating whether dropout improves performance.
+(a) Node classification tasks.
+Dropout Homophilic Networks Heterophilic Networks
+Cora CiteSeer PubMed Chameleon Squirrel TwitchDE
+DropEdge $+0.483$ $+0.828$ $-0.064$ $-1.988$ $-1.125$ $-0.206$
+DropNode $+0.200$ $-0.002$ $-0.181$ $-6.090$ $-2.016$ $-2.383$
+DropAgg $+0.322$ $+0.797$ $+0.032$ $-7.779$ $-4.904$ $-1.014$
+DropGNN $+0.519$ $+1.058$ $-0.149$ $-10.572$ $-5.214$ $-1.636$
+Dropout $+0.600$ $+0.104$ $+0.265$ $-7.891$ $-2.773$ $-1.598$
+DropMessage $+0.389$ $+0.039$ $+0.633$ $-1.569$ $-0.092$ $-0.139$
+(b) Graph classification tasks.
+Dropout Molecular Networks Social Networks
+Mutag Proteins Enzymes Reddit IMDb Collab
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+G.2 Effect Size in Statistical Tests
+Table 7: Hedges’ $g$ statistic for different dataset--model--dropout combinations.
+Color-coding for effect size according to Cohen [14]; red denotes negative effect, and green denotes positive effect.
+Medium to large positive effect sizes in bold.
+X No effect    X X Small effect    X X Medium effect    X X Large effect
+Dataset GNN DropEdge DropNode DropAgg DropGNN Dropout DropMessage
+Cora GCN $\mathbf{+0.973}$ $\mathbf{+0.759}$ $\mathbf{+0.431}$ $\mathbf{+0.900}$ $\mathbf{+0.990}$ $+0.060$
+GIN $-0.174$ $+0.205$ $+0.068$ $-0.212$ $\mathbf{+1.010}$ $\mathbf{+2.180}$
+GAT $\mathbf{+0.933}$ $+0.317$ $\mathbf{+0.455}$ $\mathbf{+1.083}$ $\mathbf{+0.977}$ $\mathbf{+0.710}$
+CiteSeer GCN $\mathbf{+0.848}$ $+0.268$ $\mathbf{+0.628}$ $\mathbf{+0.929}$ $+0.155$ $-0.397$
+GIN $-0.254$ $\mathbf{+0.515}$ $-0.124$ $-0.854$ $+0.006$ $+0.132$
+GAT $\mathbf{+0.799}$ $-0.002$ $\mathbf{+0.777}$ $\mathbf{+0.985}$ $+0.103$ $+0.040$
+PubMed GCN $\mathbf{+1.039}$ $\mathbf{+1.285}$ $-0.627$ $\mathbf{+1.455}$ $\mathbf{+2.625}$ $\mathbf{+2.907}$
+GIN $-0.043$ $+0.343$ $-0.546$ $-1.213$ $\mathbf{+0.423}$ $\mathbf{+1.118}$
+GAT $-0.108$ $-0.320$ $+0.053$ $-0.271$ $\mathbf{+0.407}$ $\mathbf{+0.964}$
+Chameleon GCN $-0.167$ $-0.174$ $-3.279$ $-0.479$ $-0.385$ $\mathbf{+0.356}$
+GIN $-0.219$ $-0.567$ $-0.329$ $-0.510$ $-0.857$ $+0.290$
+GAT $-0.429$ $-1.326$ $-1.945$ $-2.438$ $-1.504$ $-0.333$
+Squirrel GCN $+0.006$ $-0.389$ $-8.183$ $-0.180$ $-0.066$ $+0.189$
+GIN $+0.179$ $-0.204$ $+0.199$ $-0.052$ $-0.111$ $+0.085$
+GAT $-0.393$ $-0.750$ $-1.899$ $-2.251$ $-1.165$ $-0.038$
+TwitchDE GCN $-0.099$ $-0.113$ $-2.272$ $-0.426$ $-0.280$ $+0.165$
+GIN $-0.079$ $-0.116$ $+0.140$ $-0.005$ $-0.389$ $-0.016$
+GAT $-0.201$ $-1.206$ $-0.726$ $-0.938$ $-0.921$ $-0.129$
+Mutag GCN $-0.087$ $-0.448$ $-0.069$ $-0.017$ $-0.026$ $+0.174$
+GIN $-0.107$ $-0.351$ $-0.330$ $-0.640$ $-0.240$ $-0.342$
+GAT $+0.071$ $+0.034$ $+0.099$ $+0.151$ $+0.091$ $+0.281$
+Proteins GCN $+0.339$ $\mathbf{+0.450}$ $+0.260$ $+0.307$ $\mathbf{+0.397}$ $\mathbf{+0.477}$
+GIN $-0.331$ $-0.541$ $-0.352$ $-0.491$ $-0.657$ $-0.217$
+GAT $-0.076$ $-0.042$ $+0.040$ $+0.031$ $-0.136$ $-0.021$
+Enzymes GCN $-0.072$ $-0.272$ $-0.055$ $-0.359$ $-0.758$ $-0.543$
+GIN $-0.673$ $-0.070$ $-0.083$ $-0.548$ $-0.290$ $-0.048$
+GAT $+0.031$ $-0.508$ $+0.055$ $-0.380$ $-0.960$ $-0.353$
+Reddit GCN $-1.712$ $-1.333$ $-3.075$ $-2.360$ $-1.065$ $-1.527$
+GIN $-0.531$ $+0.095$ $+0.163$ $+0.240$ $\mathbf{+0.503}$ $+0.165$
+GAT $+0.209$ $\mathbf{+0.638}$ $+0.154$ $-0.054$ $\mathbf{+0.632}$ $\mathbf{+0.469}$
+IMDb GCN $+0.246$ $\mathbf{+0.513}$ $\mathbf{+0.642}$ $+0.281$ $+0.200$ $+0.240$
+GIN $-0.364$ $-0.024$ $-0.327$ $-0.932$ $-0.204$ $-0.090$
+GAT $-0.030$ $\mathbf{+0.917}$ $+0.245$ $+0.344$ $\mathbf{+0.885}$ $\mathbf{+0.650}$
+Collab GCN $-0.195$ $-0.859$ $-6.962$ $+0.140$ $-0.348$ $-0.065$
+GIN $-0.220$ $\mathbf{+0.396}$ $-0.916$ $-1.139$ $-0.071$ $\mathbf{+0.396}$
+GAT $+0.000$ $+0.000$ $+0.000$ $+0.000$ $+0.000$ $+0.000$
+The reliance on p-values as a measure of statistical significance has been widely criticized due to its limitations in conveying the magnitude of an effect.
+Although a low p-value indicates that an observed difference is unlikely to have occurred under the null hypothesis, it does not provide information about the practical significance of the result.
+A statistically significant effect may be too small to be meaningful in real-world applications, while a non-significant result does not necessarily imply the absence of a meaningful effect, particularly when sample sizes are small.
+These concerns have led to an increased emphasis on effect size measures, which quantify the magnitude of differences independently of sample size.
+One widely used measure of effect size is Cohen’s $d$ statistic [14], which standardizes the difference between two group means by dividing by the pooled standard deviation:
+$d$ $=\frac{\bar{x}_{1}-\bar{x}_{2}}{s_{p}}$ (G.1)
+$s_{p}$ $=\sqrt{\frac{\left(n_{1}-1\right)s_{1}^{2}+\left(n_{2}-1\right)s_{2}^{2}}{n_{1}+n_{2}-2}}$ (G.2)
+where $\bar{x}_{1}$ and $\bar{x}_{2}$ are sample means, $s_{1}^{2}$ and $s_{2}^{2}$ are unbiased sample variance, and $n_{1}$ and $n_{2}$ are the sizes of the samples.
+However, Cohen’s $d$ assumes that the sample standard deviation is an unbiased estimator of the population standard deviation.
+In small samples, this assumption does not hold, as the sample standard deviation tends to underestimate the true population variability.
+To address this bias, Hedges’ $g$ statistic [39] introduces a correction factor that adjusts Cohen’s $d$ statistic for small sample sizes:
+$g\approx\left(1-\frac{3}{4\left(n_{1}+n_{2}\right)-9}\right)d$ (G.3)
+Cohen [14] suggested that an effect size of $0.2$ be considered small, $0.5$ be considered medium, and $0.8$ be considered large.
+In Table 7, we present Hedges’ $g$ statistic for the statistical tests in Section 5.2.
+We can clearly see that for homophilic datasets, there is a strong positive effect of using the dropping methods, but for heterophilic datasets and graph classification datasets, there is at most a small positive effect of using the dropping methods; rather, in most cases there is a negative effect.
+G.3 Performance of GIN with DropSens
+Table 8: Performance of GIN with different rewiring methods in graph-classification tasks, following [44, 8]. First, second, and third best results are colored.
+Rewiring Mutag Proteins Enzymes Reddit IMDb Collab
+None $77.70$ $70.80$ $33.80$ $86.79$ $70.18$ 72.99
+Last FA 83.45 72.30 47.40 90.22 70.91 75.06
+Every FA $72.55$ $70.38$ $28.38$ $50.36$ $49.16$ $32.89$
+DIGL 79.70 $70.76$ 35.72 $76.04$ $64.39$ $54.50$
+SDRF 78.40 $69.81$ 35.82 $86.44$ $69.72$ $72.96$
+FoSR $78.00$ 75.11 $29.20$ 87.35 71.21 73.28
+GTR $77.60$ 73.13 $30.57$ 86.98 71.28 $72.93$
+DropSens $70.60$ $68.00$ $29.56$ $76.44$ $62.48$ $65.77$
+In Section 5.3, we showed that when modelling long-range graph-classification tasks using GCNs, DropSens outperforms state-of-the-art graph-rewiring techniques designed for alleviating over-squashing.
+However, it does not perform as well with GIN, as can be seen in Table 8 – unsurprising, since DropSens was specifically designed to work with GCN’s message-passing scheme.
+This highlights the main limitation of DropSens, necessitating architecture-specific alteration to the edge-dropping strategy, which is not practical in general.
+G.4 Best-performing Dropping Probabilities
+Table 9: Best performing dropout configuration – $q_{\max}$ and $c$ for DropSens, and $q$ for other dropping methods.
+GNN Dataset DropEdge DropNode DropAgg DropGNN Dropout DropMessage DropSens
+Cora GCN $0.7$ $0.4$ $0.1$ $0.7$ $0.7$ $0.1$ $0.5$, $0.95$
+GIN $0.1$ $0.2$ $0.1$ $0.1$ $0.3$ $0.5$ $0.2$, $0.90$
+GAT $0.8$ $0.3$ $0.9$ $0.4$ $0.7$ $0.6$ $0.8$, $0.50$
+CiteSeer GCN $0.9$ $0.1$ $0.9$ $0.8$ $0.1$ $0.3$ $0.2$, $0.95$
+GIN $0.1$ $0.4$ $0.1$ $0.1$ $0.1$ $0.2$ $0.3$, $0.95$
+GAT $0.9$ $0.2$ $0.9$ $0.8$ $0.1$ $0.1$ $0.8$, $0.50$
+PubMed GCN $0.3$ $0.4$ $0.1$ $0.2$ $0.5$ $0.7$ $0.5$, $0.95$
+GIN $0.1$ $0.1$ $0.1$ $0.1$ $0.1$ $0.6$ $0.2$, $0.90$
+GAT $0.1$ $0.1$ $0.1$ $0.1$ $0.5$ $0.8$ $0.3$, $0.95$
+Chameleon GCN $0.4$ $0.1$ $0.1$ $0.1$ $0.4$ $0.5$ $0.5$, $0.80$
+GIN $0.1$ $0.1$ $0.2$ $0.1$ $0.1$ $0.1$ $0.2$, $0.95$
+GAT $0.1$ $0.1$ $0.1$ $0.1$ $0.2$ $0.1$ $0.2$, $0.95$
+Squirrel GCN $0.2$ $0.6$ $0.9$ $0.3$ $0.5$ $0.6$ $0.5$, $0.90$
+GIN $0.2$ $0.5$ $0.2$ $0.5$ $0.3$ $0.7$ $0.2$, $0.95$
+GAT $0.1$ $0.1$ $0.1$ $0.1$ $0.1$ $0.1$ $0.2$, $0.95$
+TwitchDE GCN $0.3$ $0.2$ $0.9$ $0.1$ $0.1$ $0.6$ $0.3$, $0.95$
+GIN $0.2$ $0.3$ $0.1$ $0.2$ $0.1$ $0.5$ $0.2$, $0.90$
+GAT $0.1$ $0.1$ $0.1$ $0.1$ $0.1$ $0.1$ $0.2$, $0.90$
+Actor GCN $0.9$ $0.1$ $0.9$ $0.5$ $0.2$ $0.1$ $0.2$, $0.95$
+GIN $0.5$ $0.1$ $0.2$ $0.3$ $0.6$ $0.2$ $0.2$, $0.90$
+GAT $0.9$ $0.2$ $0.8$ $0.7$ $0.1$ $0.1$ $0.8$, $0.50$
+Mutag GCN $0.5$ $0.8$ $0.7$ $0.2$ $0.2$ $0.3$ $0.5$, $0.80$
+GIN $0.1$ $0.3$ $0.1$ $0.6$ $0.2$ $0.5$ $0.5$, $0.95$
+GAT $0.4$ $0.1$ $0.2$ $0.1$ $0.4$ $0.4$ $0.5$, $0.95$
+Proteins GCN $0.1$ $0.1$ $0.1$ $0.2$ $0.1$ $0.3$ $0.8$, $0.90$
+GIN $0.1$ $0.3$ $0.1$ $0.1$ $0.4$ $0.1$ $0.5$, $0.95$
+GAT $0.4$ $0.1$ $0.2$ $0.1$ $0.1$ $0.1$ $0.8$, $0.95$
+Enzymes GCN $0.5$ $0.1$ $0.1$ $0.6$ $0.1$ $0.4$ $0.8$, $0.80$
+GIN $0.3$ $0.1$ $0.1$ $0.1$ $0.1$ $0.1$ $0.8$, $0.95$
+GAT $0.1$ $0.1$ $0.1$ $0.1$ $0.1$ $0.1$ $0.5$, $0.90$
+Reddit GCN $0.1$ $0.1$ $0.1$ $0.1$ $0.2$ $0.1$ $0.2$, $0.95$
+GIN $0.1$ $0.1$ $0.1$ $0.1$ $0.1$ $0.1$ $0.2$, $0.95$
+GAT $0.9$ $0.7$ $0.6$ $0.7$ $0.6$ $0.9$ $0.8$, $0.95$
+IMDb GCN $0.6$ $0.7$ $0.1$ $0.6$ $0.1$ $0.9$ $0.8$, $0.50$
+GIN $0.1$ $0.3$ $0.1$ $0.1$ $0.2$ $0.2$ $0.2$, $0.95$
+GAT $0.7$ $0.7$ $0.1$ $0.6$ $0.9$ $0.6$ $0.2$, $0.95$
+Collab GCN $0.1$ $0.1$ $0.4$ $0.1$ $0.1$ $0.1$ $0.2$, $0.95$
+GIN $0.1$ $0.2$ $0.1$ $0.1$ $0.2$ $0.1$ $0.2$, $0.90$
+GAT $0.1$ $0.1$ $0.1$ $0.1$ $0.1$ $0.1$ $0.2$, $0.90$
+For the real-world datasets in Section 5.2, we report the best performing dropping probability for different dataset--model--dropout combinations in Table 9.